### v1

# Conditional-$t^{3}$VAE: Equitable Latent Space Allocation for Fair Generation

Abstract
Variational Autoencoders (VAEs) with global priors mirror the training set’s class frequency in latent space, underrepresenting tail classes and reducing generative fairness on imbalanced datasets. While $t^{3}$VAE improves robustness via heavy-tailed Student’s t-distribution priors, it still allocates latent volume proportionally to the class frequency. In this work, we address this issue by explicitly enforcing equitable latent space allocation across classes. To this end, we propose Conditional-$t^{3}$VAE, which defines a per-class Student’s t joint prior over latent and output variables, preventing dominance by majority classes. Our model is optimized using a closed-form objective derived from the $\gamma$-power divergence. Moreover, for class-balanced generation, we derive an equal-weight latent mixture of Student’s t-distributions. On SVHN-LT, CIFAR100-LT, and CelebA, Conditional-$t^{3}$VAE consistently achieves lower FID scores than both $t^{3}$VAE and Gaussian-based VAE baselines, particularly under severe class imbalance. In per-class F1 evaluations, Conditional-$t^{3}$VAE also outperforms the conditional Gaussian VAE across all highly imbalanced settings. While Gaussian-based models remain competitive under mild imbalance ratio ($\rho\lesssim 3$), our approach substantially improves generative fairness and diversity in more extreme regimes.
1 Introduction
Class imbalance and long-tail distributions are prevalent in real-world datasets, yet generative models often fail to represent rare classes accurately. When trained on skewed data, these models tend to overfit dominant modes and underrepresent minority ones in latent and output spaces, resulting in biased or unfair generations. This issue is especially critical in sensitive applications such as facial synthesis [1] and medical imaging [2], where such biases can exacerbate social and diagnostic disparities [3].
Variational Autoencoders (VAEs) [4] are a widely used class of generative models, valued for their probabilistic formulation, stable training, and compatibility with latent-variable modeling frameworks leading in image quality such as Latent Diffusion Models (LDMs) [5]. Although GANs and diffusion models often achieve lower FID scores, VAEs offer unique advantages in class-conditional generation, interpretability, and efficient inference, making them strong candidates for improving fairness under class imbalance. Standard VAEs commonly use isotropic Gaussian priors, which inadequately model heavy-tailed structures and rare phenomena [6]. Prior efforts to address this have introduced non-Gaussian priors, particularly Student’s t-distributions [7, 8, 9, 10], to enhance robustness. However, these approaches often rely on global priors, causing the latent space to be dominated by majority classes under skewed distributions.
We address this issue with Conditional-$t^{3}$VAE (C-$t^{3}$VAE), a conditional generative model that imposes a per-class Student’s t-distribution prior over the joint latent-output space. This design ensures allocating an equal latent space volume per class, thereby mitigating majority class dominance, while the heavy tails of the Student’s t-distribution more effectively capture intra-class variation. To enable class-balanced sampling, we introduce an equal-weight mixture of Student’s t-distributions with analytically derived component variances. Together, these components enable balanced class-conditional generation and mitigate bias present in unconditional models. We summarize our main theoretical and empirical contributions in the following points :
  • We propose the C-$t^{3}$VAE model with a training objective based on the $\gamma$-power divergence.
  • We develop an equal-weight latent mixture sampling scheme with analytically derived optimal variance scaling for each component.
  • We outperform relevant baselines in FID on SVHN-LT [11], CIFAR100-LT [12], and CelebA [13] under severe imbalance, and show via per-class evaluation that C-$t^{3}$VAE better avoids mode collapse, exceeding a conditional VAE in per class Recall and F1 while remaining competitive on Precision.
  • We identify the imbalance ratio threshold $\rho\approx 3$, beyond which Gaussian priors become suboptimal, providing guidance for model selection on skewed datasets.
2 Related Work
Since the introduction of VAEs [4], many extensions have sought to improve latent representation by replacing the standard Gaussian prior with more expressive alternatives. These include Gaussian mixtures [14, 15], hyperspherical priors [16], normalizing flows [17], Riemannian priors [18], and implicit distributions [19]. While most retain the Evidence Lower Bound (ELBO) formulation, others adopt alternative objectives or divergence measures for added flexibility.
To address long-tailed or imbalanced data, Student’s t-distributions have been explored for their robustness and heavy tails [6]. Methods such as [7] and [8] model the latent space of the autoencoder through a t-distributed prior, and rely on KL-divergence-based ELBO objectives. However, since the KL divergence lacks a closed-form solution for the Student’s t-distributions, these methods resort to numerical approximation. The $t^{3}$VAE [10] improves on this by modeling the joint latent-output distribution and optimizing a closed-form $\gamma$-divergence objective [9]. Nonetheless, it still employs a global latent prior, which results in a latent space volume allocation reflecting class frequency and leading to imbalance present in the generated samples.
Other works address fairness in generative modeling using normalizing flows [17] and diffusion models [20, 21]. However, since leading image generation models are based on Latent Diffusion Models [5] and VAEs are a corner-stone in the design of these models, we deem it crucial to improve VAEs’ ability to handle imbalanced data. Consequently, as fairness in latent space allocation remains under-explored in VAEs, we introduce a class-conditional, heavy-tailed prior to address this gap and allow for balanced latent space allocation across classes.
3 Background
This section introduces the theoretical background and baseline models relevant to our work. We assume access to a labeled, imbalanced dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}\in\mathbb{R}^{n}$ is a data sample of dimension $n$, $y_{i}\in\{1,\dots,K\}$ its class label and $m$ the latent space dimension.
3.1 VAEs and Conditional-VAEs
VAEs [4] are generative models trained via variational inference by maximizing the ELBO of the log-likelihood. The standard objective of this model is
$\mathcal{L}_{\theta,\phi}:=\mathbb{E}_{z\sim q_{\phi}(\cdot|x)}[\log p_{\theta}(x|z)]-\mathcal{D}_{KL}(q_{\phi}(z|x)\|p(z)),$ (1)
where the first term is the reconstruction loss with $p_{\theta}(x|z)$ being the decoder. The second term is the Kullback–Leibler (KL) divergence between the approximate posterior $q_{\phi}(z|x)$ and the prior $p(z)$. The $\beta$-VAE is a weighted variant of the VAE model which introduces a $\beta$ scaling term for the KL divergence [22]:
$\mathcal{L}_{\theta,\phi}:=\mathbb{E}_{z\sim q_{\phi}(\cdot|x)}[\log p_{\theta}(x|z)]-\beta\mathcal{D}_{KL}(q_{\phi}(z|x)\|p(z)),$ (2)
with $p(z)\sim\mathcal{N}_{m}(0,I)$, $q_{\phi}(\cdot|x)\sim\mathcal{N}_{m}(\mu_{\phi}(x),\Sigma_{\phi}(x))$, and $p_{\theta}(x|z)\sim\mathcal{N}_{m}(\mu_{\theta}(z),\sigma^{2}I)$. $\mu_{\phi}(\cdot)$ and $\Sigma_{\phi}(\cdot)$ are the mean and the covariance matrices inferred through a neural network with parameter $\phi$ given the input $x$. This variant of the VAE model allows to place more weight on disentangling the latent space or on the reconstruction of the data points. To generate samples from the VAE or the $\beta$-VAE model, we sample a latent vector $z\sim\mathcal{N}_{m}(0,I)$. Then, the generated data point would be $\hat{x}\sim\mathcal{N}_{m}(\mu_{\theta}(z),\sigma^{2}I)$.
Nevertheless, since Eq. (1) and Eq. (2) optimize the ELBO over the data distribution $p_{data}(x)$, which can be decomposed as $p_{\mathrm{data}}(x)=\sum_{y_{i}}p(y_{i})\,p_{\mathrm{data}}(x\mid y_{i}),$ this optimization inherently biases the model toward head classes with larger $p(y_{i})$. As a result, most generated samples come from overrepresented classes, while tail classes’ samples are underrepresented and of lower quality, a phenomenon commonly referred to as mode collapse. Therefore, when labels are available, it is preferable to define class-conditional approximate posterior and prior distributions: $q_{\phi}(z|x,y)$ and $p(z|y)$. This yields the Conditional-VAE (CVAE) model trained using the objective [23]:
$\sum_{y}\ \mathbb{E}_{z\sim q_{\phi}(\cdot|x,y)}[\log p_{\theta}(x|z,y)]-\beta\mathcal{D}_{KL}(q_{\phi}(z|x,y)\|p(z|y)).$ (3)
Here, in Eq. (3) we constrain all $p(y_{i})$ to be equal and omit them from the loss function in order not to exacerbate the issue of class imbalance in the latent space. Also, we define $p(z|y)\sim\mathcal{N}_{m}(\mu_{y},I)$ with learnable class-wise means $\mu_{y}$. To generate a data point $\hat{x}_{y}$ from class $y$, we sample $z_{y}\sim\mathcal{N}_{m}(\mu_{y},I)$, then we get $\hat{x}_{y}\sim p_{\theta}(x|z_{y},y)$. Nevertheless, despite conditioning, this formulation remains Gaussian. Unlike Student’s t-distributions, Gaussian priors poorly approximate heavy-tailed data distributions [6].
3.2 Multivariate Student’s t-Distribution and $\gamma$-Power Divergence
A $d$-dimensional Student’s t-distribution with mean $\mu\in\mathbb{R}^{d}$, covariance $\Sigma\in\mathbb{R}^{d\times d}$, and degrees of freedom $\nu>2$ is a heavy-tail, super-Gaussian distribution defined as
$t_{d}(x)=C_{\nu,d}|\Sigma|^{-\frac{1}{2}}\left(1+\frac{(x-\mu)^{\top}\Sigma^{-1}(x-\mu)}{\nu}\right)^{-\frac{\nu+d}{2}},\qquad C_{\nu,d}=\frac{\Gamma\left(\frac{\nu+d}{2}\right)}{\Gamma\left(\frac{\nu}{2}\right)\left(\nu\pi\right)^{\frac{d}{2}}}.$ (4)
The power form of this distribution prevents a closed-form KL divergence between two Student’s t-distributions. Instead, the $\gamma$-power divergence $\mathcal{D}_{\gamma}(q\|p)$ is used [9, 10]. This divergence is defined for $q\sim t_{d}(\mu_{0};\Sigma_{0};\nu)$, $p\sim t_{d}(\mu_{1};\Sigma_{1};\nu)$ starting from the $\gamma$-entropy $\mathcal{H}_{\gamma}(p)$ and cross-entropy $\mathcal{C}_{\gamma}(q,p)$
$\mathcal{H}_{\gamma}(p):=-\|p\|_{1+\gamma}$ $=-\left(\int p(x)^{1+\gamma}dx\right)^{\frac{1}{1+\gamma}},\qquad\mathcal{C}_{\gamma}(q,p):=-\int q(x)\left(\frac{p(x)}{\|p\|_{1+\gamma}}\right)^{\gamma}dx.$
$\mathcal{D}_{\gamma}(q\|p):=\gamma^{-1}\left(\mathcal{C}_{\gamma}(q,p)-\mathcal{H}_{\gamma}(p)\right)$ (5)
with $\gamma=-\frac{2}{\nu+d}$. Then, substituting the definition of a Student’s t-distribution from Eq. (4) into Eq. (5), the following closed-form formula for the $\gamma$-power divergence can be derived (See Appendix A):
$\mathcal{D}_{\gamma}(q\|p)$ $=-\frac{C_{\nu,d}^{\frac{\gamma}{1+\gamma}}}{\gamma}\left(1+\frac{d}{\nu-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg{[}-|\Sigma_{0}|^{-\frac{\gamma}{2(1+\gamma)}}\quad\left(1+\frac{d}{\nu-2}\right)$
$\quad+|\Sigma_{1}|^{-\frac{\gamma}{2}}|\Sigma_{0}|^{\frac{\gamma^{2}}{2(1+\gamma)}}\Bigg{(}1+\frac{\operatorname{Tr}\left(\Sigma_{1}^{-1}\Sigma_{0}\right)}{\nu-2}+\frac{(\mu_{0}-\mu_{1})^{\top}\Sigma_{1}^{-1}(\mu_{0}-\mu_{1})}{\nu}\Bigg{)}\Bigg{]}.$ (6)
3.3 $t^{3}$-Variational Autoencoder
3.3.1 Definition
The $t^{3}$VAE model [10] is a non-ELBO-based autoencoder which models the joint prior distribution $p_{\theta}(x,z)$ using multivariate Student’s t-distributions
$p_{\theta}(x,z)\propto\sigma^{-n}\left[1+\frac{1}{\nu}\left(\|z\|^{2}+\frac{\|x-\mu_{\theta}(z)\|^{2}}{\sigma^{2}}\right)\right]^{-\frac{\nu+m+n}{2}},$
where $\mu_{\theta}(\cdot)$ is the decoder neural network with parameter $\theta$ whereas $\sigma$ is a parameter controlling the decoder’s output covariance. From this joint distribution, the marginal latent prior $p(z)$ and decoder distribution $p_{\theta}(x|z)$ can be defined. Furthermore, the approximate posterior distribution is defined as :
$q_{\phi}(z|x)=t_{m}\left(x\middle|\mu_{\phi}(x),\frac{\Sigma_{\phi}(x)}{1+\nu^{-1}n},\nu+n\right).$
Hence, the data distribution would be $q_{\phi}(x,z)=p_{\text{data}}(x)q_{\phi}(z|x)$. As a result, relying on the $\gamma$-divergence in Eq. (6) applied to the $p_{\theta}(x,z)$ and $q_{\phi}(x,z)$ distributions, the following loss function is derived to optimize the $t^{3}$VAE’s parameters :
$\mathcal{L}_{\gamma}=\mathbb{E}_{x}\Bigg{[}\frac{\mathbb{E}_{z}\left[\|x-\mu_{\theta}(z)\|^{2}\right]}{\sigma^{2}}+\|\mu_{\phi}(x)\|^{2}+\frac{\nu\operatorname{Tr}\left(\Sigma_{\phi}(x)\right)}{\nu+n-2}-\frac{\nu C_{1}}{C_{2}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}\Bigg{]},$ (7)
with $\gamma=-\frac{2}{\nu+n+m}$ and $C_{1}$ and $C_{2}$ being theoretically derived constants. We note that the first term in this loss function represents the standard reconstruction term in VAE models and the rest of the terms are regularization terms over the latent space. To sample from the latent space of the $t^{3}$VAE, [10] propose the $p_{\nu}^{\star}(z)=t_{m}(0,\tau^{2}I,\nu+n)$ distribution with
$\tau^{2}=\frac{1}{1+\nu^{-1}n}\left(\frac{C_{\nu,n}}{\sigma^{n}}\cdot\frac{\nu-2}{\nu+n-2}\right)^{\frac{2}{\nu+n-2}}.$ (8)
Moreover, sampling from a multi-variate Student’s t-distribution $T\sim t_{d}(\mu,\Sigma,\nu)$ both in the learning (Eq. (7)) and sampling (Eq. (8)) phases is performed through the standard reparameteration trick for Student’s t-distributions $T:=\mu+Z\sqrt{\nu V^{-1}}$ where $Z\sim\mathcal{N}(0,\Sigma)$ and $V\sim\mathcal{X}^{2}(\nu)$.
3.3.2 $\beta$-$t^{3}$VAE
From Eq. (7) we can also define a $\beta$-$t^{3}$VAE model by multiplying all the regularization terms by a $\beta$ factor. Similarly to $\beta$-VAE models, this improves the versatility of the model and allows either a focus on generation or disentangling.
3.3.3 $\tau^{2}$ Improvement
Closely analyzing the proposed derivation of $\tau^{2}$ of the $t^{3}$VAE model, we discovered a subtle issue in its mathematical formulation. The employed $\gamma$-power divergence presents a discrepancy to the correct formula in Eq. (6). We revised the formulation and the corrected $\tau^{2}$ is (See Appendix D):
$\tau^{2}$ $=\frac{1}{1+\nu^{-1}n}\left(|\Sigma_{\phi}(x)|^{\frac{\gamma}{2}}\frac{C_{\nu,n}}{\sigma^{n}}\cdot\frac{\nu-2}{\nu+n-2}\right)^{\frac{2\gamma}{(1+\gamma)(2+\gamma m)}}\approx\frac{1}{1+\nu^{-1}n}\left(\frac{C_{\nu,n}}{\sigma^{n}}\cdot\frac{\nu-2}{\nu+n-2}\right)^{-\frac{2\gamma}{2+\gamma m}}.$ (9)
The corrected exact form of $\tau^{2}$ is applicable when $|\Sigma_{\phi}(x)|$ is known and when the dimension of the data is low. However, for high dimensional data, as handled in this work, we get $\gamma\approx 0$. Hence, one can use the approximation without any loss in accuracy. We note that the new form of $\tau^{2}$ leads to a similar empirical value of standard deviation compared to the previous form. Nevertheless, for correctness, in our sampling from the latent space of the $t^{3}$VAE model we use the approximation in Eq. (9).
In summary, although the $t^{3}$VAE effectively models heavy-tailed distributions through Student’s t-distributions and $\gamma$-power divergence, it does not explicitly address class imbalance in the latent space by not allocating equal volume for each class. In the next section, we introduce a class-conditional variant of the $t^{3}$VAE, designed to ensure fair and balanced generation across all classes.
4 Conditional $t^{3}$-Variational Autoencoder
We propose the Conditional $t^{3}$-Variational Autoencoder (C-$t^{3}$VAE), present its formulation, training objective, and sampling strategy. C-$t^{3}$VAE models the latent space as a mixture of Student’s t-distributions, one per class, ensuring equal latent volume allocation and promoting fairness in generation. Intra-class imbalance is further handled by the heavy-tailed nature of the Student’s t-distribution prior.
4.1 Model definition
The C-$t^{3}$VAE we propose is based on the following class conditional joint prior distribution
$p_{\theta}(x,z|y)=\frac{C_{\nu,m+n}}{|\Sigma_{x}|^{\frac{1}{2}}|\Sigma_{y}|^{\frac{1}{2}}}\Bigg{[}1+\frac{(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})+(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))}{\nu}\Bigg{]}^{-\frac{\nu+m+n}{2}},$
with $\nu$, $n$ and $m$ being the degrees of freedom of the Student’s t-distribution, the dimension of the input data and the dimension of the latent space respectively. $\mu_{y}\in\mathbb{R}^{m}$ is a learnable mean vector representing class centers in latent space of dimension $m$. Moreover, $\Sigma_{x}$ and $\Sigma_{y}$ are the covariance matrices of the prior distributions over the latent and output variables.
From this joint distribution, we can derive the conditional latent prior $p(z|y)=t_{m}(z|\mu_{y},\Sigma_{y},\nu)$ and decoder distribution (See Appendix B)
$p_{\theta}(x|z,y)=t_{n}\Bigg{(}x\Bigg{|}\mu_{\theta}(z),\frac{\left(1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right)}{(1+\nu^{-1}m)}\Sigma_{x},\nu+m\Bigg{)}.$
Furthermore, as in $t^{3}$VAE, we define the approximate posterior $q_{\phi}(z|x)$ as a multivariate Student’s t-distribution capturing heavy-tailed structure in the latent space :
$q_{\phi}(z|x)=t_{m}\left(z\middle|\mu_{\phi}(x),\frac{\Sigma_{\phi}(x)}{1+\nu^{-1}n},\nu+n\right).$
4.2 Objective function
Harnessing Eq. (5) and the defined prior and posterior distributions of the proposed C-$t^{3}$VAE, we derive in Appendix C the following class-wise objective
$\mathcal{L}(\gamma,y)$ $=\mathbb{E}_{x}\Bigg{[}\mathbb{E}_{z}\Big{[}(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))\Big{]}+(\mu_{\phi}(x)-\mu_{y})^{\top}\Sigma_{y}^{-1}(\mu_{\phi}(x)-\mu_{y})$
$\quad+\frac{\nu\operatorname{Tr}\left(\Sigma_{y}^{-1}\Sigma_{\phi}(x)\right)}{\nu+n-2}-\frac{\nu C_{1}}{C_{2}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}\Bigg{]},$
with $C_{1}=\left(C^{\gamma}_{\nu+n,m}\left(1+\frac{n}{\nu}\right)^{\frac{\gamma m}{2}}\frac{\nu+n+m-2}{\nu+n-2}\right)^{\frac{1}{1+\gamma}}$ and $C_{2}=\left(\frac{C_{\nu,m+n}^{\gamma}}{|\Sigma_{x}|^{\frac{\gamma}{2}}|\Sigma_{y}|^{\frac{2\gamma+1}{2}}}\left(1+\frac{m+n}{\nu-2}\right)^{-\gamma}\right)^{\frac{1}{1+\gamma}}$. By taking $\Sigma_{x}=\sigma^{2}I$ and $\Sigma_{y}=I$, $\mathcal{L}(\gamma,y)$ objective function simplifies to :
$\mathcal{L}(\gamma,y)=\mathbb{E}_{x}\Bigg{[}\frac{\mathbb{E}_{z}\left[\|x-\mu_{\theta}(z)\|^{2}\right]}{\sigma^{2}}+\|\mu_{\phi}(x)-\mu_{y}\|^{2}+\frac{\nu\operatorname{Tr}\left(\Sigma_{\phi}(x)\right)}{\nu+n-2}-\frac{\nu C_{1}}{C_{2}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}\Bigg{]}.$ (10)
Therefore, we express the final loss function $\mathcal{L}(\gamma)$ over the whole dataset as : $\mathcal{L}(\gamma)=\sum_{y}\mathcal{L}(\gamma,y).$ As in Eq. (3), here too we consider all $p(y_{i})$ to be equal and hence rule out their contribution to the loss function. This is done to avoid emphasizing the imbalance present in the data.
4.3 Sampling distribution
Similarly to the objective function of $t^{3}$VAE, $\mathcal{L}(\gamma,y)$ in Eq. (10) can be decomposed into a reconstruction term and regularization terms. To sample from the latent space of the C-$t^{3}$VAE, we focus on the regularization terms and define the following sampling distribution :
$p_{\nu}^{\star}(z)=\sum_{y=1}^{K}\alpha_{y}\cdot t_{m}(\mu_{y},\tau^{2}I,\nu+n),\quad\forall y,\;\;\alpha_{y}=\frac{1}{K}.$ (11)
The theoretical derivation of the variance $\tau^{2}$ leads to the form expressed in Eq. (9) (Derivation in Appendix D).
The mixture-based sampling distribution we define in Eq. (11) with equal $\alpha_{y}$ ensures a uniformly sampled synthetic data across all classes, regardless of their frequency in the original training data. As a result, C-$t^{3}$VAE equipped with this sampling distribution mitigates the common problem in generative models where head-class samples dominate due to their density in latent space. Furthermore, by modifying the mixture weights $\alpha_{y}$, one can prioritize specific classes. This makes our method flexible for targeted data augmentation or fairness-aware sampling strategies.
4.3.1 $\beta$-C-$t^{3}$VAE
As with $t^{3}$VAE, the class-wise objective defined in Eq. (10) can be split into a reconstruction and regularization terms. By preceding the regularization term with a $\beta$ scalar, we can define a $\beta$-C-$t^{3}$VAE model thereby improving the domain of applicability of the model.
Overall, C-$t^{3}$VAE provides a principled, flexible, and tractable framework for fair generative modeling, particularly under class-imbalanced conditions. In the following, we study the C-$t^{3}$VAE’s performance across multiple datasets with varying imbalance degrees.
5 Experiments
In this section we outline the generative performance of the proposed C-$t^{3}$VAE model on labeled datasets compared to relevant VAE baselines [^1] We restrict our comparisons to VAE-based generative models, as Latent Diffusion Models [5], the state-of-the-art in image generation, depend on VAEs as a core component. Enhancing fairness in VAEs under class imbalance is therefore a necessary precursor to advancing more complex models.. We conduct experiments on three datasets notably SVHN-LT [11], CIFAR100-LT [24, 12] and CelebA [13] each chosen to highlight different challenges.
5.1 Evaluation procedure
All models are evaluated using Fréchet Inception Distance (FID) [25], computed against a balanced test set for each dataset. This setup measures how effectively a model overcomes training set imbalance by assessing its ability to generate high-quality samples across all classes. To evaluate robustness, we impose varying degrees of imbalance during training.
For SVHN-LT and CIFAR100-LT, we introduce class imbalance by applying an exponential decay to the number of samples per class after equalizing class sizes in the original train and test sets. The imbalance ratio $\rho$ defines the ratio between the most and least frequent classes, with class-wise sample counts $M_{y_{i}}$ given by: $M_{y_{i}}=M\cdot\rho^{-\frac{y_{i}-1}{K-1}},$ where $M$ is the original per-class sample count.
For CelebA, we compute FID per attribute, treating each attribute (eg. Mustache) and its negation (eg. no Mustache) as separate binary classes. The training set uses CelebA’s inherent imbalance, while the test set is balanced by downsampling to the smaller class size. Multi-attribute generation is not considered in this work and is left for future exploration.
5.2 Results
We present both quantitative and qualitative results of the C-$t^{3}$VAE model and the models it directly improves upon notably the VAE, C-VAE and $t^{3}$VAE models with their $\beta$ variants. This controlled comparison helps isolate the contributions of key design choices:
  • VAE : ELBO trained standard Gaussian-based VAEs.
  • C-VAE : VAE supplemented by conditional Gaussian priors to assess the class conditioning effect without changing the prior family.
  • $t^{3}$VAE : Student’s t-distribution latent prior and $\gamma$-power divergence objective, does not use class-conditional priors, nor allows for class conditional generation. Through a comparison with this model we assess the role of conditional modeling on Student’s t-distribution priors.
In the following, we analyze the latent sampling standard deviation $\tau$ employed in $t^{3}$VAE and C-$t^{3}$VAE models, varying it to assess alignment between empirical and theoretical values. Then, we present FID comparisons across baselines with optimized hyper-parameters (The tuning of $\beta$, $\nu$ and $\tau$ hyper-parameters is reported in Appendix F), followed by per-class generative evaluation.
(a) SVHN-LT
(b) CIFAR100-LT
(c) CelebA
Figure 1: FID score as a function of $\tau$ for the $t^{3}$VAE and C-$t^{3}$VAE models. Results are for the imbalance ratio $\rho=100$ for the SVHN-LT and CIFAR100-LT, and for the Mustache attribute ($\rho=25$) in the case of the CelebA dataset. Other imbalance ratios’ results paint a similar picture and are provided in Appendix F.3. The horizontal dashed lines is the FID value of the best performing VAE and C-VAE on each dataset and the vertical dashed line is the value of $\tau$ as derived in Eq. (9). We note that the used models in these figures have optimized $\beta$ and $\nu$ hyper-parameters.
5.2.1 $\tau$ parameter study
Figure 1 illustrates the sampling standard deviation’s $\tau$ impact on the FID metric. We observe that models based on the Student’s t-distribution benefit from higher standard deviation on CIFAR100-LT compared to SVHN-LT. Specifically, C-$t^{3}$VAE outperforms the C-VAE for $\tau\in[0.25;0.55]$ on CIFAR100-LT and $\tau\in[0.19;0.28]$ on SVHN-LT and for all $\tau$ values on the CelebA dataset. Moreover, it surpasses the $t^{3}$VAE models’ FID for all $\tau$ values and across all datasets.
For both Student’s t-distribution based models, the optimal FID score for the SVHN-LT occurs near the theoretically derived $\tau$ value. However, for the more complex CIFAR100-LT dataset, the optimal $\tau$ is higher than the theoretical value $\tau=0.4$. We hypothesize that the higher variance reflects the model’s need during training to accommodate the greater complexity and entropy of the dataset, which makes it harder to align the learned approximate posterior with the prior and to produce confident reconstructions. For the CelebA dataset, the standard deviation parameter $\tau$ has minimal to no impact on model performance likely due to the lower variability in the dataset’s images as they all represent faces.
5.2.2 Optimized Model Results Discussion
After optimizing the hyperparameters of the various models tested in this work, we present their generation FID scores in Table 1. We provide results for $\beta$ models after optimization and non $\beta$ models ($\beta=1$) to underscore the importance of this parameter on the resulting performance as it was not explored in the $t^{3}$VAE work [10].
Table 1: Generation FID results on the SVHN-LT, CIFAR100-LT and CelebA datasets. For the SVHN-LT and CIFAR100-LT datasets we use different imbalance ratios $\rho\in\{100,50,10,1\}$. However, for the CelebA dataset we use the Mustache, Young, Male and Smiling attributes which have imbalance ratios of $25$, $3.5$, $1.4$ and 11 respectively. The $\beta$ models undertook an optimization of the $\beta$ hyper-parameter while non-$\beta$ models have $\beta=1$. All models have optimized $\nu$ and $\tau$ hyper-parameters. The attributes for the CelebA dataset column indicate which attribute is used to condition the conditional models and balance the test set.
SVHN-LT CIFAR100-LT CelebA
Models $\rho=$100 50 10 1 $\rho=$100 50 10 1 Mustache Young Male Smiling
VAE 93.89 91.91 91.66 92.16 163.66 162.91 165.47 166.46 110.58 92.01 110.58 82.05
$\beta$-VAE 47.11 49.81 45.70 43.48 122.62 123.07 123.72 124.43 104.63 92.87 087.96 83.15
C-VAE 74.75 70.40 72.30 74.16 157.90 163.67 162.09 163.24 096.98 89.17 086.17 78.35
$\beta$-C-VAE 48.39 46.39 43.97 43.87 114.88 118.89 114.89 118.21 098.35 85.53 079.76 78.46
$t^{3}$VAE 57.07 54.30 52.10 51.52 136.63 137.24 138.92 135.23 105.80 88.07 083.62 78.90
$\beta$-$t^{3}$VAE 51.62 49.55 48.93 45.37 109.11 107.93 108.97 111.00 105.86 88.21 083.83 78.89
C-$t^{3}$VAE 47.09 46.29 47.43 51.32 125.48 127.96 130.28 129.40 101.18 87.07 081.92 80.97
$\beta$-C-$t^{3}$VAE 44.02 42.60 42.01 44.49 103.25 102.99 105.92 112.37 095.82 82.61 081.65 80.08
Figure 2: Sample synthetic images from the optimized VAE and $t^{3}$VAE models trained on the CelebA dataset. No class conditioning is possible for these models.
From Table 1, $t^{3}$VAE consistently improves over the VAE model. This underlines the generative advantage of the Student’s t-distribution prior over a Gaussian one, in addition to the reconstruction advantage highlighted by [10]. Moreover, optimizing the $\beta$ hyper-parameter improves $t^{3}$VAE’s FID over the VAE on the CIFAR100-LT and CelebA, while remaining competitive on SVHN-LT. Also, qualitative results of Figure 2 show that the $t^{3}$VAE model compared to the VAE model produces significantly sharper synthetic images on the CelebA dataset.
Figure 3: Sample synthetic images for the optimized C-VAE and C-$t^{3}$VAE trained on specific attributes of the CelebA dataset.
For class-conditional models, $\beta$ optimized C-$t^{3}$VAE yields strong FID improvements across all imbalance settings, surpassing all baselines. As per Table 1, it achieves a gain of up to 4, 5 and 10 FID points over the $\beta$-$t^{3}$VAE model on the imbalanced settings of the SHVN-LT, CIFAR100-LT and CelebA, respectively. This underlines the importance of an equal per-class volume in an imbalanced dataset setting. Moreover, $\beta$-C-$t^{3}$VAE reduces the FID by up to 4 and 15 points over C-VAE on SVHN-LT and CIFAR100-LT respectively as per Table 1. For the CelebA dataset, Table 1 shows that $\beta$-C-$t^{3}$VAE achieves the best results on heavily imbalanced attributes like Mustache, demonstrating better generation for underrepresented groups. This gain in performance compared to the C-VAE goes back to the use of Student’s t-distribution latent prior and its ability to better capture intra-class long-tail distributions. Additionally, qualitative samples of conditional optimized models (Figure 3) reveal sharper facial features in C-$t^{3}$VAE compared to the C-VAE. These results establish it as the most reliable model across imbalance ratios and resolutions.
In summary, C-$t^{3}$VAE after optimization of all hyper-parameters notably $\beta$, $\nu$ and $\tau$ consistently achieves the lowest FID, validating the use of class-conditional Student’s t-distribution priors and balanced latent sampling. Improvements on CIFAR100-LT are especially pronounced, highlighting the model’s robustness to severe imbalance. However, for balanced settings, C-$t^{3}$VAE remains competitive.
5.2.3 Per-class evaluation
In this section, we evaluate the conditional models on a per-class basis. However, since FID is biased on small datasets and offers limited insight as a single scalar metric, we rely on Precision, Recall, and F1 generative metrics [^2] Precision quantifies sample quality (sharpness), Recall reflects mode coverage and F1 is their harmornic mean. Together they allows us to gain deeper insight into the models’ behavior. [26]. Our results on CelebA are shown in Figure 4, with additional results for SVHN-LT and CIFAR100-LT included in Appendix G.
(a) Recall
(b) Precision
(c) F1 score
Figure 4: Per-class generative metrics on CelebA after optimization of all hyper-parameters notably $\beta$, $\nu$ and $\tau$. We note that the imbalance ratio of the Mustache, Young, Male and Smiling factors $\rho$ are $25$, $3.5$, $1.4$ and 11 respectively.
Figure 5: Fine-grained comparison of C-$t^{3}$VAE and C-VAE models under varying imbalance ratios on SVHN-LT.
On the CelebA dataset (Figure 4), C-$t^{3}$VAE improves Recall and F1 on the most imbalanced attribute (Mustache), but not on more balanced ones (Male, Smiling) which is an expected behavior. However, surprisingly, for the attribute Young (imbalance ratio 3.5), C-VAE performs slightly better this suggests that there is a regime where Gaussian priors may still suffice. To explore this, we vary the imbalance ratio on SVHN-LT from $100$ to $1$ and plot the results in Figure 5. This figure shows there is a threshold $\rho\approx 3$ before which C-VAE performs better; beyond that, C-$t^{3}$VAE has the advantage. Also, we notice that the more the imbalance ratio increases, the more FID gap becomes more significant and in favor of the C-$t^{3}$VAE.
For the SVHN-LT dataset (Figures in Appendix G), when comparing the C-$t^{3}$VAE to the C-VAE, the C-$t^{3}$VAE achieves significantly higher Recall and competitive Precision. This indicates a better coverage of data modes and maintaining images quality which yields higher F1 scores especially on tail classes. For the CIFAR100-LT dataset (Figures in Appendix G), C-VAE often achieves high Precision but zero Recall, indicating mode collapse. In contrast, the C-$t^{3}$VAE preserves Recall at the cost of slightly lower Precision, also resulting in improved F1 scores. Consequently, on the SVHN-LT CIFAR100-LT and CelebA datasets, C-$t^{3}$VAE sets itself as the most reliable method for fair, high quality image generation.
6 Conclusion
We introduced C-$t^{3}$VAE, a class-conditional generative model that leverages Student’s t-distributions in the latent space with a theoretically derived balanced sampling scheme. This approach improves fairness and sample quality under class imbalance. Extensive experiments on SVHN-LT, CIFAR100-LT, and CelebA demonstrate that, after the optimization of the $\beta$, $\nu$ and $\tau$ hyper-parameters, C-$t^{3}$VAE consistently outperforms the $t^{3}$VAE and conditional VAE baselines, achieving up to a 15-point FID improvement on highly imbalanced datasets. Per-class Precision, Recall, and F1 evaluations confirm improved mode coverage, particularly in tail classes. Moreover, we identify $\rho>3$ as the regime where C-$t^{3}$VAE outperforms Gaussian-prior models, with the performance gap widening as $\rho$ increases. For $\rho\lesssim 3$, C-$t^{3}$VAE remains competitive with C-VAE.
Beyond generative performance, C-$t^{3}$VAE contributes to fairness-aware modeling by promoting balanced latent sampling and better mode coverage in underrepresented classes. Future work will explore extending C-$t^{3}$VAE to multi-label problems and integration with Latent Diffusion models.
Acknowledgement
The authors were granted access to the HPC resources of MesoPSL financed by the Région Île-de-France and the Equip@Meso project (reference ANR-10-EQPX-29-01) of the programme investissements d’avenir supervised by France’s Agence nationale pour la recherche.
Appendix for
”Conditional-$t^{3}$VAE: Equitable Latent Space Allocation for Fair Generation”
Appendix A $\gamma$-power divergence corrected derivation
In this section, we present our correction of the previously derived form of the $\gamma$-power divergence in [10]. We start from the integrals derived in [10] and keep a similar notation :
$\int q(x)p(x)^{\gamma}dx$ $=C_{\nu,d}^{\gamma}|\Sigma_{1}|^{-\frac{\gamma}{2}}\left(1+\frac{1}{\nu-2}\operatorname{Tr}(\Sigma_{1}^{-1}\Sigma_{0})+\frac{1}{\nu}(\mu_{0}-\mu_{1})^{\top}\Sigma_{1}^{-1}(\mu_{0}-\mu_{1})\right)$
$\int q(x)^{1+\gamma}dx$ $=C_{\nu,d}^{\gamma}|\Sigma_{0}|^{-\frac{\gamma}{2}}\left(1+\frac{d}{\nu-2}\right)$
$\int p(x)^{1+\gamma}dx$ $=C_{\nu,d}^{\gamma}|\Sigma_{1}|^{-\frac{\gamma}{2}}\left(1+\frac{d}{\nu-2}\right)$
Combining these formulas gives a similar enthropy $\mathcal{H}_{\gamma}$ to the one derived in [10]:
$\mathcal{H}_{\gamma}(q)=-C_{\nu,d}^{\frac{\gamma}{1+\gamma}}|\Sigma_{0}|^{-\frac{\gamma}{2(1+\gamma)}}\left(1+\frac{d}{\nu-2}\right)^{\frac{1}{1+\gamma}}.$
However, the cross-entropy $\mathcal{C}_{\gamma}$ takes the following form :
[…]
In red, we highlight the main difference to the formula derived in [10]. $\mathcal{H}_{\gamma}(q)$ and $\mathcal{C}_{\gamma}(q,p)$ combine to give :
$\mathcal{D}_{\gamma}(q\|p)$ $=-\frac{C_{\nu,d}^{\frac{\gamma}{1+\gamma}}}{\gamma}\left(1+\frac{d}{\nu-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg{[}-|\Sigma_{0}|^{-\frac{\gamma}{2(1+\gamma)}}\left(1+\frac{d}{\nu-2}\right)$
[…]
Appendix B Priors derivations
In this section, we present our derivations of the different prior distributions defining our proposed C-$t^{3}$-VAE model. Starting from the proposed joint distribution :
$p_{\theta}(x,z|y)=\frac{C_{\nu,m+n}}{|\Sigma_{x}|^{\frac{1}{2}}|\Sigma_{y}|^{\frac{1}{2}}}\left[1+\frac{(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})+(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))}{\nu}\right]^{-\frac{\nu+m+n}{2}}.$
To calculate the prior distribution on the latent space we marginalize out $x$ as follows :
$p(z|y)$ $=\int p_{\theta}(x,z|y)dx$
$=\int C_{\nu,m+n}|\Sigma_{x}|^{-\frac{1}{2}}|\Sigma_{y}|^{-\frac{1}{2}}\Bigg{[}1+\frac{(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})}{\nu}$
$\quad+\frac{(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))}{\nu}\Bigg{]}^{-\frac{\nu+m+n}{2}}dx$
$=C_{\nu,m+n}|\Sigma_{x}|^{-\frac{1}{2}}|\Sigma_{y}|^{-\frac{1}{2}}\left[1+\frac{1}{\nu}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right]^{-\frac{\nu+m+n}{2}}$
$\times\int\left(1+\frac{(1+\nu^{-1}m)(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))}{\left(1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right)(\nu+m)}\right)^{-\frac{\nu+m+n}{2}}dx.$
Given that :
$\int C_{\nu+m,n}|\Sigma|^{-\frac{1}{2}}\left(1+\frac{(x-\mu)^{\top}\Sigma^{-1}(x-\mu)}{\nu+m}\right)^{-\frac{\nu+m+n}{2}}dx$ $=1$
$\Rightarrow\int\left(1+\frac{(x-\mu)^{\top}\Sigma^{-1}(x-\mu)}{\nu+m}\right)^{-\frac{\nu+m+n}{2}}dx$ $=C_{\nu+m,n}^{-1}|\Sigma|^{\frac{1}{2}},$
and when setting :
$\Sigma^{-1}=\frac{(1+\nu^{-1}m)\Sigma_{x}^{-1}}{1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})},$
We get :
$\int\left(1+\frac{(1+\nu^{-1}m)(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))}{\left(1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right)(\nu+m)}\right)^{-\frac{\nu+m+n}{2}}dx$
$\qquad\qquad\qquad\qquad\qquad\qquad=C_{\nu+m,n}^{-1}\left|\left(\frac{(1+\nu^{-1}m)\Sigma_{x}^{-1}}{1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})}\right)^{-1}\right|^{\frac{1}{2}}$
$\qquad\qquad\qquad\qquad\qquad\qquad=C_{\nu+m,n}^{-1}\left|\frac{1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})}{(1+\nu^{-1}m)}\Sigma_{x}\right|^{\frac{1}{2}}$
$\qquad\qquad\qquad\qquad\qquad\qquad=C_{\nu+m,n}^{-1}\left(\frac{1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})}{1+\nu^{-1}m}\right)^{\frac{n}{2}}\left|\Sigma_{x}\right|^{\frac{1}{2}}.$
Therefore, $p(z|y)$ simplifies to :
$p(z|y)$ $=C_{\nu,m+n}|\Sigma_{x}|^{-\frac{1}{2}}|\Sigma_{y}|^{-\frac{1}{2}}\left[1+\frac{1}{\nu}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right]^{-\frac{\nu+m+n}{2}}C_{\nu+m,n}^{-1}\left|\Sigma_{x}\right|^{\frac{1}{2}}$
$\quad\times\left(\frac{1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})}{1+\nu^{-1}m}\right)^{\frac{n}{2}}$
$=C_{\nu,m+n}C_{\nu+m,n}^{-1}\left(1+\frac{m}{\nu}\right)^{-\frac{n}{2}}|\Sigma_{y}|^{-\frac{1}{2}}\left[1+\frac{1}{\nu}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right]^{-\frac{\nu+m}{2}}$
$=C_{\nu,m}|\Sigma_{y}|^{-\frac{1}{2}}\left[1+\frac{1}{\nu}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right]^{-\frac{\nu+m}{2}}$
$=t_{m}(z|\mu_{y},\Sigma_{y},\nu).$
Here and in the following, we use the fact
$C_{\nu,m+n}=C_{\nu+m,n}C_{\nu,m}\left(1+\frac{m}{\nu}\right)^{\frac{n}{2}}.$
Besides, the prior distribution over the output of the decoder model $p(x|z,y)$ can be derived as follows :
$p_{\theta}(x|z,y)$ $=\frac{p_{\theta}(x,z|y)}{p(z|y)}$
$=\frac{C_{\nu,m+n}}{|\Sigma_{x}|^{\frac{1}{2}}|\Sigma_{y}|^{\frac{1}{2}}}\left[1+\frac{(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})+(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))}{\nu}\right]^{-\frac{\nu+m+n}{2}}$
$\quad\times C_{\nu,m}^{-1}|\Sigma_{y}|^{\frac{1}{2}}\left[1+\frac{1}{\nu}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right]^{\frac{\nu+m}{2}}$
$=C_{\nu+m,n}|\Sigma_{x}|^{-\frac{1}{2}}\left(1+\frac{m}{\nu}\right)^{\frac{n}{2}}\left[1+\frac{1}{\nu}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right]^{-\frac{n}{2}}$
$\quad\times\left(1+\frac{(1+\nu^{-1}m)(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))}{\left(1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right)(\nu+m)}\right)^{-\frac{\nu+m+n}{2}}$
$=t_{n}\left(x\middle|\mu_{\theta}(z),\frac{\left(1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right)}{(1+\nu^{-1}m)}\Sigma_{x},\nu+m\right).$
Appendix C Loss function derivation
In this section, we derive the loss function of C-$t^{3}$-VAE. We start by calculating the different double integrals $\iint p_{\theta}(x,z|y)^{1+\gamma}dxdz$, $\iint q_{\phi}(x,z|y)p_{\theta}(x,z|y)^{\gamma}dxdz$, and $\iint q_{\phi}(x,z|y)^{1+\gamma}dxdz$.
Firstly,
$\iint p_{\theta}(x,z|y)^{1+\gamma}dxdz$ $=\mathbb{E}_{z\sim p(z|y)}\mathbb{E}_{x\sim p_{\theta}(x|z,y)}\left[p_{\theta}(x,z|y)^{\gamma}\right]$
$=C_{\nu,m+n}^{\gamma}|\Sigma_{x}|^{-\frac{\gamma}{2}}|\Sigma_{y}|^{-\frac{\gamma}{2}}\mathbb{E}_{z}\mathbb{E}_{x}\Bigg{[}1+\frac{(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})}{\nu}$
$\quad+\frac{(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))}{\nu}\Bigg{]}$
$=C_{\nu,m+n}^{\gamma}|\Sigma_{x}|^{-\frac{\gamma}{2}}|\Sigma_{y}|^{-\frac{\gamma}{2}}\mathbb{E}_{z}\bigg{[}1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})$
$\quad+\nu^{-1}\mathbb{E}_{x}\left[\operatorname{Tr}(\Sigma_{x}^{-1}(x-\mu_{\theta}(z))(x-\mu_{\theta}(z))^{\top})\right]\bigg{]}$
$=C_{\nu,m+n}^{\gamma}|\Sigma_{x}|^{-\frac{\gamma}{2}}|\Sigma_{y}|^{-\frac{\gamma}{2}}\mathbb{E}_{z}\Bigg{[}1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})$
$\quad+\nu^{-1}\operatorname{Tr}\left(\Sigma_{x}^{-1}\Sigma_{x}\frac{\nu+m}{\nu+m-2}\frac{\left(1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right)}{(1+\nu^{-1}m)}\right)\Bigg{]}$
Here, we use the following identities
$(k-p)^{\top}H^{-1}(k-p)=\operatorname{Tr}\left(H^{-1}(k-p)(k-p)^{\top}\right);\qquad\mathbb{E}[\operatorname{Tr}(\cdot)]=\operatorname{Tr}(\mathbb{E}[\cdot])$
and the covariance of a multivariate Student’s t distribution $p\sim t(\mu;\Sigma;\nu)$ is $\frac{\nu}{\nu-2}\Sigma$. Consequently, and after a few simplifications we get
$\iint p_{\theta}(x,z|y)^{1+\gamma}dxdz$ $=C_{\nu,m+n}^{\gamma}|\Sigma_{x}|^{-\frac{\gamma}{2}}|\Sigma_{y}|^{-\frac{\gamma}{2}}\mathbb{E}_{z}\bigg{[}1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})$
$\quad+\frac{n}{\nu+m-2}\left(1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right)\bigg{]}$
$=C_{\nu,m+n}^{\gamma}|\Sigma_{x}|^{-\frac{\gamma}{2}}|\Sigma_{y}|^{-\frac{\gamma}{2}}\mathbb{E}_{z}\Bigg{[}\left(1+\frac{n}{\nu+m-2}\right)$
$\quad\times\Big{(}1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\Big{)}\Bigg{]}$
$=C_{\nu,m+n}^{\gamma}|\Sigma_{x}|^{-\frac{\gamma}{2}}|\Sigma_{y}|^{-\frac{\gamma}{2}}\left(1+\frac{n}{\nu+m-2}\right)$
$\quad\times\Big{(}1+\nu^{-1}\mathbb{E}_{z}\left[(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right]\Big{)}$
$=C_{\nu,m+n}^{\gamma}|\Sigma_{x}|^{-\frac{\gamma}{2}}|\Sigma_{y}|^{-\frac{\gamma}{2}}\left(1+\frac{n}{\nu+m-2}\right)$
$\quad\times\Big{(}1+\nu^{-1}\mathbb{E}_{z}\left[\operatorname{Tr}(\Sigma_{y}^{-1}(z-\mu_{y})(z-\mu_{y})^{\top}\right]\Big{)}$
$=C_{\nu,m+n}^{\gamma}|\Sigma_{x}|^{-\frac{\gamma}{2}}|\Sigma_{y}|^{-\frac{\gamma}{2}}\left(1+\frac{n}{\nu+m-2}\right)\left(1+\frac{m}{\nu-2}\right)$
$=C_{\nu,m+n}^{\gamma}|\Sigma_{x}|^{-\frac{\gamma}{2}}|\Sigma_{y}|^{-\frac{\gamma}{2}}\left(1+\frac{m+n}{\nu-2}\right).$
Secondly,
$\iint q_{\phi}(x,z|y)p_{\theta}(x,z|y)^{\gamma}dxdz$ $=\mathbb{E}_{x\sim p_{data}}\mathbb{E}_{z\sim q(z|x)}\left[p_{\theta}(x,z|y)^{\gamma}\right]$
$=C_{\nu,m+n}^{\gamma}|\Sigma_{x}|^{-\frac{\gamma}{2}}|\Sigma_{y}|^{-\frac{\gamma}{2}}\mathbb{E}_{x}\mathbb{E}_{z}\Bigg{[}1+\frac{(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})}{\nu}$
$\quad+\frac{(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))}{\nu}\Bigg{]}$
$=C_{\nu,m+n}^{\gamma}|\Sigma_{x}|^{-\frac{\gamma}{2}}|\Sigma_{y}|^{-\frac{\gamma}{2}}\mathbb{E}_{x}\Bigg{[}1+\frac{1}{\nu}\mathbb{E}_{z}\left[(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right]$
$\quad+\frac{1}{\nu}\mathbb{E}_{z}\left[(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))\right]\Bigg{]}.$
Simplifying $\mathbb{E}_{z}\left[(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right]$ :
$\mathbb{E}_{z}\left[(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right]$ $=\mathbb{E}_{z}\left[\operatorname{Tr}\left(\Sigma_{y}^{-1}(z-\mu_{y})(z-\mu_{y})^{\top}\right)\right]$
$=\mathbb{E}_{z}\left[\operatorname{Tr}\left(\Sigma_{y}^{-1}(z-\mu(x)+\mu(x)-\mu_{y})(z-\mu(x)+\mu(x)-\mu_{y})^{\top}\right)\right]$
$=\mathbb{E}_{z}[\operatorname{Tr}(\Sigma_{y}^{-1}((z-\mu(x))(z-\mu(x))^{\top}+(z-\mu(x))(\mu(x)-\mu_{y})^{\top}$
$+(\mu(x)-\mu_{y})(z-\mu(x))^{\top}+(\mu(x)-\mu_{y})(\mu(x)-\mu_{y})^{\top}))]$
$=\frac{\nu}{\nu+n-2}\operatorname{Tr}\left(\Sigma_{y}^{-1}\Sigma_{\phi}(x)\right)+(\mu(x)-\mu_{y})^{\top}\Sigma_{y}^{-1}(\mu(x)-\mu_{y}).$
Then, $\iint q(x,z|y)p_{\theta}(x,z|y)^{\gamma}dxdz$ simplifies to :
$\iint q_{\phi}(x,z|y)p_{\theta}(x,z|y)^{\gamma}dxdz$ $=C_{\nu,m+n}^{\gamma}|\Sigma_{x}|^{-\frac{\gamma}{2}}|\Sigma_{y}|^{-\frac{\gamma}{2}}\mathbb{E}_{x}\Bigg{[}1+\frac{1}{\nu}\mathbb{E}_{z}\left[(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right]$
$\quad+\frac{1}{\nu}\mathbb{E}_{z}\left[(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))\right]\Bigg{]}$
$\iint q_{\phi}(x,z|y)p_{\theta}(x,z|y)^{\gamma}dxdz$ $=C_{\nu,m+n}^{\gamma}|\Sigma_{x}|^{-\frac{\gamma}{2}}|\Sigma_{y}|^{-\frac{\gamma}{2}}\,|\Sigma_{y}|^{-\frac{1}{2}}\,\mathbb{E}_{x}\Bigg{[}1+\frac{1}{\nu}\frac{\nu\operatorname{Tr}\left(\Sigma_{y}^{-1}\Sigma_{\phi}(x)\right)}{\nu+n-2}$
$\quad+\frac{(\mu(x)-\mu_{y})^{\top}\Sigma_{y}^{-1}(\mu(x)-\mu_{y})}{\nu}$
$\quad+\frac{\mathbb{E}_{z}\left[(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))\right]}{\nu}\Bigg{]}.$
Finally, the third term $\iint q_{(}x,z|y)^{1+\gamma}dxdz$ is
$\iint q_{\phi}(x,z|y)^{1+\gamma}dxdz=C^{\gamma}_{\nu+n,m}\left(1+\frac{n}{\nu}\right)^{\frac{\gamma m}{2}}\left(1+\frac{m}{\nu+n-2}\right)\int|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2}}p_{data}(x)^{1+\gamma}dx,$
where this last double integral is equal to the one computed for the $t^{3}$-VAE.
Equipped with these formulas we can calculate the entropy $\mathcal{H}_{\gamma}$, cross-entropy $\mathcal{C}_{\gamma}$ and the $\gamma$-divergence $\mathcal{D}(q||p)$ of our model. Firstly,
$\mathcal{H}_{\gamma}$ $=-\left(\iint q(x,z)^{1+\gamma}dxdz\right)^{\frac{1}{1+\gamma}}$
$=-C^{\frac{\gamma}{1+\gamma}}_{\nu+n,m}\left(1+\frac{n}{\nu}\right)^{\frac{\gamma m}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)^{\frac{1}{1+\gamma}}\left(\int|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2}}p_{data}(x)^{1+\gamma}dx\right)^{\frac{1}{1+\gamma}},$
Which is similar to the one calculated in the $t^{3}$VAE model.
Secondly,
$\mathcal{C}_{\gamma}$ $=-\left(\iint q(x,z|y)p_{\theta}(x,z|y)^{\gamma}dxdz\right)\left(\iint p_{\theta}(x,z|y)^{1+\gamma}\right)^{-\frac{\gamma}{1+\gamma}}$
$=-C_{\nu,m+n}^{\gamma}|\Sigma_{x}|^{-\frac{\gamma}{2}}|\Sigma_{y}|^{-\frac{\gamma}{2}-\frac{1}{2}}\mathbb{E}_{x}\Bigg{[}1+\frac{1}{\nu}\frac{\nu\operatorname{Tr}\left(\Sigma_{y}^{-1}\Sigma_{\phi}(x)\right)}{\nu+n-2}+\frac{(\mu(x)-\mu_{y})^{\top}\Sigma_{y}^{-1}(\mu(x)-\mu_{y})}{\nu}$
$\quad+\frac{1}{\nu}\mathbb{E}_{z}\left[(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))\right]\Bigg{]}\left(C_{\nu,m+n}^{\gamma}|\Sigma_{x}|^{-\frac{\gamma}{2}}|\Sigma_{y}|^{-\frac{\gamma}{2}}\left(1+\frac{m+n}{\nu-2}\right)\right)^{-\frac{\gamma}{1+\gamma}}$
$=-\left(C_{\nu,m+n}^{\gamma}|\Sigma_{x}|^{-\frac{\gamma}{2}}|\Sigma_{y}|^{-\frac{\gamma}{2}}\right)^{\frac{1}{1+\gamma}}|\Sigma_{y}|^{-\frac{1}{2}}\left(1+\frac{m+n}{\nu-2}\right)^{-\frac{\gamma}{1+\gamma}}\,\mathbb{E}_{x}\Bigg{[}1+\frac{1}{\nu}\Bigg{(}\frac{\nu\operatorname{Tr}\left(\Sigma_{y}^{-1}\Sigma_{\phi}(x)\right)}{\nu+n-2}$
$\quad+(\mu(x)-\mu_{y})^{\top}\Sigma_{y}^{-1}(\mu(x)-\mu_{y})+\mathbb{E}_{z}\left[(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))\right]\Bigg{)}\Bigg{]}.$
Hence, we can define our divergence as :
$\mathcal{D}_{\gamma}(q||p)$ $=\frac{C_{1}}{\gamma}\left(\int|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2}}p_{data}(x|y)^{1+\gamma}dx\right)^{\frac{1}{1+\gamma}}-\frac{C_{2}}{\gamma}\mathbb{E}_{x}\Bigg{[}1+\frac{1}{\nu}\Bigg{(}\frac{\nu\operatorname{Tr}\left(\Sigma_{y}^{-1}\Sigma_{\phi}(x)\right)}{\nu+n-2}$
$\quad+(\mu(x)-\mu_{y})(\mu(x)-\mu_{y})^{\top}+\mathbb{E}_{z}\left[(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))\right]\Bigg{)}\Bigg{]}$
$=\mathbb{E}_{x}\Bigg{[}\frac{C_{1}}{\gamma}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}-\frac{C_{2}}{\gamma}\Bigg{(}1+\frac{1}{\nu}\Bigg{(}\frac{\nu}{\nu+n-2}\operatorname{Tr}\left(\Sigma_{y}^{-1}\Sigma_{\phi}(x)\right)$
$\quad+(\mu(x)-\mu_{y})^{\top}\Sigma_{y}^{-1}(\mu(x)-\mu_{y})+\mathbb{E}_{z}\left[(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))\right]\Bigg{)}\Bigg{)},\Bigg{]}$
with $C_{1}$ and $C_{2}$ being :
$C_{1}$ $=C^{\frac{\gamma}{1+\gamma}}_{\nu+n,m}\left(1+\frac{n}{\nu}\right)^{\frac{\gamma m}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)^{\frac{1}{1+\gamma}}$
$C_{2}$ $=\left(C_{\nu,m+n}^{\gamma}|\Sigma_{x}|^{-\frac{\gamma}{2}}|\Sigma_{y}|^{-\frac{2\gamma+1}{2}}\right)^{\frac{1}{1+\gamma}}\left(1+\frac{m+n}{\nu-2}\right)^{-\frac{\gamma}{1+\gamma}}.$
On that account, the loss function for a class $y$ is :
$\mathcal{L}(\gamma,y)$ $=-\frac{\nu\gamma}{C_{2}}\cdot\mathcal{D}_{\gamma}(q||p)$
$=\mathbb{E}_{x}\Bigg{[}\mathbb{E}_{z}\left[(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))\right]+(\mu(x)-\mu_{y})^{\top}\Sigma_{y}^{-1}(\mu(x)-\mu_{y})$
$\quad+\frac{\nu}{\nu+n-2}\operatorname{Tr}\left(\Sigma_{y}^{-1}\Sigma_{\phi}(x)\right)-\frac{\nu C_{1}}{C_{2}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}\Bigg{]},$
and by taking $\Sigma_{x}=\sigma^{2}I$ and $\Sigma_{y}=I$, we obtain :
[…]
Appendix D Sampling distribution variance derivation
In this section, we present the derivation of $\tau^{2}$ used in the sampling of $t^{3}$VAE and C-$t^{3}$VAE model. We present only the derivation for the C-$t^{3}$-VAE and it is identical to the one for the $t^{3}$-VAE since the former model is a generalization of the later.
First, we simplify the divergence $\mathcal{D}(q\|p^{\star})$ :
$\mathcal{D}(q\|p^{\star})$ $=-\frac{C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}}{\gamma}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg{[}-\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}$
$\quad\times\left(1+\frac{m}{\nu+n-2}\right)+|\tau^{2}I|^{-\frac{\gamma}{2}}|\left(1+\nu^{-1}n\right)^{-1}\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}$
$\quad\times\left(1+\frac{\operatorname{Tr}\left(\tau^{-2}\left(1+\nu^{-1}n\right)^{-1}\Sigma_{\phi}(x)\right)}{\nu+n-2}+\frac{\tau^{-2}}{\nu+n}\|\mu(x)-\mu_{y})\|^{2}\right)\Bigg{]}$
Here, we use the fact that $|\alpha A|^{\delta}=\alpha^{\delta n}|A|^{\delta}$ where $n$ is the dimension of the square $A$ matrix. Also, we use $\operatorname{Tr}(\alpha A)=\alpha\operatorname{Tr}(A)$. After simplification and rearranging we get :
$\mathcal{D}(q\|p^{\star})$ $=-\frac{1}{\gamma}C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg{[}-\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}$
$\quad\times\left(1+\frac{m}{\nu+n-2}\right)+\tau^{-\gamma m}\left(1+\nu^{-1}n\right)^{-\frac{\gamma^{2}m}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}$
$\quad\left(1+\frac{\tau^{-2}\left(1+\nu^{-1}n\right)^{-1}}{\nu+n-2}\operatorname{Tr}\left(\Sigma_{\phi}(x)\right)+\frac{\tau^{-2}}{\nu+n}\|\mu(x)-\mu_{y})\|^{2}\right)\Bigg{]}$
$=-\frac{1}{\gamma}C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg{[}-\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}$
$\quad\times\left(1+\frac{m}{\nu+n-2}\right)+\nu^{-1}\tau^{-2-\gamma m}\left(1+\nu^{-1}n\right)^{-\frac{\gamma^{2}m}{2(1+\gamma)}-1}|\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}$
$\quad\left(\kappa+\frac{\nu}{\nu+n-2}\operatorname{Tr}\left(\Sigma_{\phi}(x)\right)+\|\mu(x)-\mu_{y})\|^{2}\right)\Bigg{]}$
$=-\frac{1}{\gamma}C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\nu^{-1}\tau^{-2-\gamma m}\left(1+\nu^{-1}n\right)^{-\frac{\gamma^{2}m}{2(1+\gamma)}-1}|\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}$
$\quad\times\Bigg{[}-\nu\tau^{2+\gamma m}\left(1+\nu^{-1}n\right)^{\frac{\gamma^{2}m+\gamma m}{2(1+\gamma)}+1}|\Sigma_{\phi}(x)|^{-\frac{\gamma^{2}}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)$
$\quad+\kappa+\frac{\nu}{\nu+n-2}\operatorname{Tr}\left(\Sigma_{\phi}(x)\right)+\|\mu(x)-\mu_{y})\|^{2}\Bigg{]},$
with:
$\kappa=\nu\tau^{2}\left(1+\nu^{-1}n\right).$
Then, we match the result to the loss function in Eq. (10) to get :
$\tau^{2+\gamma m}\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2}+1}|\Sigma_{\phi}(x)|^{-\frac{\gamma^{2}}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)=\frac{C_{1}}{C_{2}}.$
Moreover, we have :
$\frac{C_{1}}{C_{2}}$ $=C^{\frac{\gamma}{1+\gamma}}_{\nu+n,m}\left(1+\frac{n}{\nu}\right)^{\frac{\gamma m}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)^{\frac{1}{1+\gamma}}C_{\nu,m+n}^{-\frac{\gamma}{1+\gamma}}\sigma^{\frac{n\gamma}{1+\gamma}}\left(1+\frac{m+n}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
$=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{\gamma}{1+\gamma}}_{\nu+n,m}C_{\nu,m+n}^{-\frac{\gamma}{1+\gamma}}\left(1+\frac{n}{\nu}\right)^{\frac{\gamma m}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)^{\frac{1}{1+\gamma}}\left(1+\frac{m+n}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
$=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{-\gamma}{1+\gamma}}_{\nu,n}\left(1+\frac{m}{\nu+n-2}\right)^{\frac{1}{1+\gamma}}\left(1+\frac{m+n}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}.$
Consequently we obtain :
$\tau^{2+\gamma m}$ $\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2}+1}|\Sigma_{\phi}(x)|^{-\frac{\gamma^{2}}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\;\,=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{-\gamma}{1+\gamma}}_{\nu,n}\left(1+\frac{m}{\nu+n-2}\right)^{\frac{1}{1+\gamma}}\left(1+\frac{m+n}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
$\tau^{2+\gamma m}$ $=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{-\gamma}{1+\gamma}}_{\nu,n}\left(1+\nu^{-1}n\right)^{\frac{-\gamma m}{2}-1}|\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)^{-1}\left(1+\frac{m}{\nu+n-2}\right)^{\frac{1}{1+\gamma}}$
$\quad\times\left(1+\frac{m+n}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
$=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{-\gamma}{1+\gamma}}_{\nu,n}\left(1+\nu^{-1}n\right)^{-\frac{\gamma m}{2}-1}|\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\left(1+\frac{m+n}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
$=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{-\gamma}{1+\gamma}}_{\nu,n}\left(1+\nu^{-1}n\right)^{-\frac{\gamma m}{2}-1}|\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}\left(\frac{\nu+n-2}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
$=\left(1+\nu^{-1}n\right)^{-\frac{\gamma m}{2}-1}\left(\sigma^{n}C^{-1}_{\nu,n}|\Sigma_{\phi}(x)|^{\frac{\gamma}{2}}\quad\frac{\nu+n-2}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}.$
Hence, we get :
$\tau^{2}$ $=\left(1+\nu^{-1}n\right)^{-\frac{\gamma m+2}{\gamma m+2}}\left(\sigma^{n}C^{-1}_{\nu,n}|\Sigma_{\phi}(x)|^{\frac{\gamma}{2}}\frac{\nu+n-2}{\nu-2}\right)^{\frac{2\gamma}{(1+\gamma)(2+\gamma m)}}$
$=\left(1+\nu^{-1}n\right)^{-1}\left(\sigma^{n}C^{-1}_{\nu,n}|\Sigma_{\phi}(x)|^{\frac{\gamma}{2}}\frac{\nu+n-2}{\nu-2}\right)^{\frac{2\gamma}{(1+\gamma)(2+\gamma m)}},$
which is the form of $\tau^{2}$ we report in Eq. (9).
Appendix E Experimental Setup Details
E.1 Datasets
We conduct experiments on three datasets notably SVHN-LT [11], CIFAR100-LT [24, 12] and CelebA [13] each chosen to highlight different challenges related to generative modeling under class imbalance and varying visual complexity.
  • SVHN-LT : This dataset is comprised of colored images of digits from 0 to 9 of size $32\times 32\times 3$. It serves as our controlled experimental setting. While simple enough for all models to converge, it is rich enough to reflect performance differences. However, as this dataset is naturally imbalanced, we balance the number of images across classes to have more control over the imposed imbalance ratio. In Table 2 we provide the number of images present in the dataset before balancing.
    Table 2: The Number of images in the SVHN dataset for the train and test sets before balancing. The value in bold is the one used to balance the dataset.
    Class 0 1 2 3 4 5 6 7 8 9
    Train set 4948 13861 10585 8497 7458 6882 5727 5595 5045 4656
    Test set 1744 5099 4149 2882 2523 2384 1977 2019 1660 1595
  • CIFAR100-LT : This balanced dataset comprised of colored images of 100 classes of different natural objects. The images are of size $32\times 32\times 3$ presents high variability with 100 fine-grained classes. In the most extreme imposed imbalance setting, tail classes may contain as few as five examples. Though the images remain low-resolution, this setting stresses the models’ robustness in the face of sparse data and various categories.
  • CelebA : This dataset enables evaluation in real-world class-imbalance scenarios using attribute labels (e.g., Mustache, Young). It also introduces the complexity of higher resolution images ($178\times 218\times 3$), testing the scalability of our models. We select four binary attributes (Mustache, Young, Male, Smiling) for evaluation, with test sets balanced by down-sampling the larger class. In Table 3, we provide the number of images for each chosen attribute in this dataset.
    Table 3: The Number of images in the CelebA dataset for the train and test sets for the Mustache, Young, Male and Smiling attributes.
    Mustache No Mustache Young Old Male Female Smiling Not Smiling
    Train set 6642 156128 126788 35982 68261 94509 78080 84690
    Test set 722 19190 15114 4848 7715 12247 9987 9975
E.2 Data Preprocessing
  • SVHN-LT Dataset For both training and testing, we crop each class to the minimum number of samples available across all classes. The only data augmentation applied is a random horizontal flip with 50% probability.
  • CIFAR100-LT Dataset We use the dataset in its entirety without class filtering. As with SVHN-LT, we apply a random horizontal flip with 50% probability for data augmentation.
  • CelebA Dataset We use the full dataset and apply a center crop of $160\times 160$ to each image and then resize to $128\times 128$. A random horizontal flip with 50% probability is also applied.
E.3 Model Architecture
Our encoder-decoder models follow a modular block design. Each encoder block consists of a convolutional layer, followed by 2D batch normalization and ReLU activation. Decoder blocks mirror this structure but replace convolutional layers with transposed convolutions.
  • SVHN-LT and CIFAR100-LT : Encoders consist of four convolutional blocks with channels {64, 128, 256, 512}, followed by two linear layers for estimating mean and covariance. The decoder uses three transposed convolutional blocks with channel sizes {128, 64, 32}, ending with a three-channel convolution and Sigmoid activation.
  • CelebA : The CelebA encoder includes six convolutional blocks with channels {64, 128, 256, 512, 512, 512}, ending with two linear layers. The decoder has six transposed convolutional layers with channels {512, 512, 256, 128, 64, 32}, followed by a final convolutional layer and Sigmoid activation.
E.4 Training Details
All models are trained using the AdamW optimizer with a learning rate of $10^{-3}$ for 150 epochs. We use a batch size of 64 for SVHN-LT and CIFAR100-LT, and 128 for CelebA.
Appendix F Hyperparameter Tuning
We present the hyperparameter tuning process used across all evaluated models. We first optimize $\beta$, then $\nu$, and finally $\tau$, yielding the models’ results reported in Table 1.
F.1 $\beta$ Optimization
We perform a hyperparameter study over $\beta$ for all tested models. Unless otherwise noted, we use the theoretically derived $\tau^{2}$ and set $\nu=10$.
F.1.1 On the SVHN-LT dataset
As shown in Figure 7, the optimal $\beta$ values for Student’s t models lies in the range $\beta\in[0.4,0.6]$ whereas it lies in the $\beta\in[0.05,0.07]$ range for Gaussian-based models. This is because the regularization term in the $\gamma$-power divergence loss is ten times larger than the KL divergence. Figure 7 also shows that FID performance is highly sensitive to $\beta$ in the Gaussian setting, requiring careful tuning which is not the case for Student’s t based models. Finally, C-$t^{3}$VAE achieves the best FID surpassing the $t^{3}$VAE and the C-VAE for all imbalance settings.
(a) SVHN-LT $\rho$=100
(b) SVHN-LT $\rho$=50
(a) SVHN-LT $\rho$=10
(b) SVHN-LT $\rho$=1
Figure 7: Variability of the FID as a function of the $\beta$ hyperparameter for the VAE, C-VAE, $t^{3}$VAE and C-$t^{3}$VAE on the SVHN-LT dataset.
F.1.2 On the CIFAR100-LT dataset
From Figure 9, we observe that Student’s t models obtain the best performance in terms of FID at $\beta=0.2$ for CIFAR100-LT dataset. However, for the Gaussian-based models, the optimal value is much lower with $\beta\in[0.02,0.05]$. The reason for this is that on this dataset too the KL regularization term is ten times smaller than the regularization terms present in the $\gamma$-power divergence loss. Additionally, we notice that C-VAE performs slightly better, likely due to the complexity of the dataset preventing full convergence to the imposed latent distribution. We further investigate this hypothesis in the $\tau$ analysis below.
(a) CIFAR100-LT $\rho$=100
(b) CIFAR100-LT $\rho$=50
(a) CIFAR100-LT $\rho$=10
(b) CIFAR100-LT $\rho$=1
Figure 9: Variability of the FID as a function of the $\beta$ hyperparameter for the VAE, C-VAE, $t^{3}$VAE and C-$t^{3}$VAE on the CIFAR100-LT dataset.
F.1.3 On the CelebA dataset
For the CelebA dataset, we focus only on $\beta$ optimization for Student’s t models. As for the Gaussian-based models, we set $\beta=0.1$, because as shown in Table 1 the optimization of the $\beta$ hyperparameter has marginal impact on the FID for the CelebA dataset, which is not the case for the SVHN-LT and CIFAR100-LT dataset. Hence, we did not deem necessary to perform a hyper-parameter optimization over $\beta$ for the Gaussian based models. Nevertheless, from Figure 11 we see that $\beta$ has marginal effect on the Student’s t based models too, likely due to the low intra-class variability.
(a) CelebA - Mustache
(b) CelebA - Young
(a) CelebA - Male
(b) CelebA - Smiling
Figure 11: Variability of the FID as a function of the $\beta$ hyperparameter for the $t^{3}$VAE and C-$t^{3}$VAE on the CelebA dataset. The horizontal lines for the VAE and C-VAE models are for the best performing model between $\beta=0.1$ and $\beta=1$.
F.2 $\nu$ Optimization
Table 4 shows results from tuning the degrees of freedom parameter $\nu$ in C-$t^{3}$VAE on both SVHN-LT and CIFAR100-LT across all imbalance ratios, using the optimal $\beta$ from the previous study. On average, $\nu=10$ performs well, consistent with prior work [10]. However, performance can be further improved by selecting $\nu$ within the range $[2.5,20]$. Still, no major influence of hyper-parameter $\nu$ on the generative FID can be observed similarly to what was observed by [10] for the reconstruction FID.
SVHN-LT CIFAR100-LT
$\nu$ 100 50 10 1 100 50 10 1
2.1 45.50 44.51 42.96 46.23 121.28 122.03 121.93 123.41
2.5 45.76 43.96 45.81 45.40 119.15 120.19 120.10 124.83
5 44.89 42.60 45.03 46.33 120.52 123.21 124.29 123.71
10 44.59 44.37 43.48 44.49 119.83 120.65 122.96 123.95
20 44.02 43.89 42.01 44.75 121.48 118.41 124.58 126.13
50 48.03 46.39 43.59 45.57 119.58 126.36 124.38 127.48
100 45.97 44.63 43.74 47.52 123.26 122.90 127.42 125.67
Table 4: Variability of the FID as a function of the standard deviation $\nu$ for the C-$t^{3}$VAE model on the SVHN-LT and CIFAR100-LT datasets.
F.3 $\tau$ Optimization
In this section, we evaluate the effect of the $\tau$ parameter on the SVHN-LT, CIFAR100-LT and CelebA datasets for all imbalance ratios while setting $\beta$ and $\nu$ to their previously optimized values. As shown in Figure 13, the optimal $\tau$ for SVHN-LT aligns closely with our theoretical prediction. In contrast, CIFAR100-LT consistently benefits from a larger $\tau=0.4$, yielding improved FID across all imbalance settings and outperforming C-VAE. On CelebA, $\tau$ has minimal impact and the most likely value is $\tau\approx 0.3$. This value is between the value observed in for the SVHN-LT and CIFAR100-LT, as is mostly liked to the complexity and the entropy of the dataset where the more complex the data is the higher the value of $\tau$ is required.
(a) SVHN-LT $\rho$=50
(b) SVHN-LT $\rho$=10
(c) SVHN-LT $\rho$=1
(d) CIFAR100-LT $\rho$=50
(e) CIFAR100-LT $\rho$=10
(f) CIFAR100-LT $\rho$=1
(a) CelebA Young
(b) CelebA Male
(c) CelebA Smiling
Figure 13: Variability of the FID as a function of the standard deviation $\tau^{2}$ for the $t^{3}$VAE and C-$t^{3}$VAE. In horizontal dashed lines is the FID value of the best performing VAE and C-VAE on each dataset. In vertical dashed lines is the theoretically identified value of $\tau$.
Appendix G Per-Class Evaluation
In this section, we assess the conditional models’ per-class Recall, Precision, and F1 metrics under all imbalance settings and for all tested datasets after optimization of all hyper-parameters.
From the following figures in Table LABEL:precision_recall_f1_shvn and LABEL:precision_recall_f1_cifar100, we see that the C-$t^{3}$VAE consistently improves Recall and mode coverage in highly imbalanced settings with $\rho=100$ and $\rho=50$. This comes at a minor Precision cost but results in significantly better F1 scores across most classes. However, on balanced or mildly imbalanced datasets, its performance remains competitive with Gaussian-based models. This observation is valid for both the SVHN-LT and CIFAR100-LT but is more pronounced on the later.
Table 5: Per-class generative metrics on SVHN-LT after optimization of $\beta$, $\nu$ and $\tau$ hyper-parameters.
$\rho$ Recall Precision F1 score
$100$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$50$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$10$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
11 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 6: Per-class generative metrics on CIFAR100-LT after optimization of $\beta$, $\nu$ and $\tau$ hyper-parameters, we focus on the top 5 head and tail classes.
$\rho$ Recall Precision F1 score
$100$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$50$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$10$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
11 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]

### v2

@@
-Conditional-$t^{3}$VAE: Equitable Latent Space Allocation for Fair Generation
+Heavy-Tailed Class-Conditional Priors for Long-Tailed Generative Modeling
 Abstract
-Variational Autoencoders (VAEs) with global priors mirror the training set’s class frequency in latent space, underrepresenting tail classes and reducing generative fairness on imbalanced datasets.
-While $t^{3}$VAE improves robustness via heavy-tailed Student’s t-distribution priors, it still allocates latent volume proportionally to the class frequency.
-In this work, we address this issue by explicitly enforcing equitable latent space allocation across classes.
-To this end, we propose Conditional-$t^{3}$VAE, which defines a per-class Student’s t joint prior over latent and output variables, preventing dominance by majority classes.
-Our model is optimized using a closed-form objective derived from the $\gamma$-power divergence.
-Moreover, for class-balanced generation, we derive an equal-weight latent mixture of Student’s t-distributions.
-On SVHN-LT, CIFAR100-LT, and CelebA, Conditional-$t^{3}$VAE consistently achieves lower FID scores than both $t^{3}$VAE and Gaussian-based VAE baselines, particularly under severe class imbalance.
-In per-class F1 evaluations, Conditional-$t^{3}$VAE also outperforms the conditional Gaussian VAE across all highly imbalanced settings.
-While Gaussian-based models remain competitive under mild imbalance ratio ($\rho\lesssim 3$), our approach substantially improves generative fairness and diversity in more extreme regimes.
+Variational Autoencoders (VAEs) with global priors trained under an imbalanced empirical class distribution can lead to underrepresentation of tail classes in the latent space.
+While $t^{3}$VAE improves robustness via heavy-tailed Student’s $t$-distribution priors, its single global prior still allocates mass proportionally to class frequency.
+We address this latent geometric bias by introducing C-$t^{3}$VAE, which assigns a per-class Student’s $t$ joint prior over latent and output variables.
+This design promotes uniform prior mass across class-conditioned components.
+To optimize our model we derive a closed-form objective from the $\gamma$-power divergence, and we introduce an equal-weight latent mixture for class-balanced generation.
+On SVHN-LT, CIFAR100-LT, and CelebA datasets, C-$t^{3}$VAE consistently attains lower FID scores than $t^{3}$VAE and Gaussian-based VAE baselines under severe class imbalance while remaining competitive in balanced or mildly imbalanced settings.
+In per-class F1 evaluations, our model outperforms the conditional Gaussian VAE across highly imbalanced settings.
+Moreover, we identify the mild imbalance threshold $\rho<5$, for which Gaussian-based models remain competitive.
+However, for $\rho\geq 5$ our approach yields improved class-balanced generation and mode coverage.
 1 Introduction
-Class imbalance and long-tail distributions are prevalent in real-world datasets, yet generative models often fail to represent rare classes accurately.
-When trained on skewed data, these models tend to overfit dominant modes and underrepresent minority ones in latent and output spaces, resulting in biased or unfair generations.
-This issue is especially critical in sensitive applications such as facial synthesis [1] and medical imaging [2], where such biases can exacerbate social and diagnostic disparities [3].
-Variational Autoencoders (VAEs) [4] are a widely used class of generative models, valued for their probabilistic formulation, stable training, and compatibility with latent-variable modeling frameworks leading in image quality such as Latent Diffusion Models (LDMs) [5].
-Although GANs and diffusion models often achieve lower FID scores, VAEs offer unique advantages in class-conditional generation, interpretability, and efficient inference, making them strong candidates for improving fairness under class imbalance.
-Standard VAEs commonly use isotropic Gaussian priors, which inadequately model heavy-tailed structures and rare phenomena [6].
-Prior efforts to address this have introduced non-Gaussian priors, particularly Student’s t-distributions [7, 8, 9, 10], to enhance robustness.
-However, these approaches often rely on global priors, causing the latent space to be dominated by majority classes under skewed distributions.
-We address this issue with Conditional-$t^{3}$VAE (C-$t^{3}$VAE), a conditional generative model that imposes a per-class Student’s t-distribution prior over the joint latent-output space.
-This design ensures allocating an equal latent space volume per class, thereby mitigating majority class dominance, while the heavy tails of the Student’s t-distribution more effectively capture intra-class variation.
-To enable class-balanced sampling, we introduce an equal-weight mixture of Student’s t-distributions with analytically derived component variances.
-Together, these components enable balanced class-conditional generation and mitigate bias present in unconditional models.
-We summarize our main theoretical and empirical contributions in the following points :
+Class imbalance and long-tail distributions are common in real-world datasets, yet generative models often fail to represent rare classes accurately.
+Under skewed training data, models tend to overfit dominant modes and underrepresent minority ones in latent and output spaces, resulting in biased generations.
+This issue is particularly consequential in applications such as facial synthesis (Mehta et al., 2024) and medical imaging (Pinaya et al., 2022), where such biases can exacerbate social and diagnostic disparities (Naik & Nushi, 2023).
+Addressing class imbalance to ensure balanced representational capacity across categories remains a major challenge for all generative models.
+For Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), their inherently unstable training dynamics, exacerbated by data imbalance, often lead to biased and mode-collapsed generations.
+Several works have sought to mitigate these effects.
+The Wasserstein GAN (Arjovsky et al., 2017) replaces the Jensen–Shannon divergence with the Wasserstein distance to stabilize optimization and improve sample diversity.
+PacGAN (Lin et al., 2018) enhances robustness by packing multiple samples into the discriminator, reducing mode collapse and improving diversity under implicit imbalances.
+Similarly, (Asokan & Seelamantula, 2020) introduces negative data augmentation to prevent under-representation of minority classes, enabling more class-balanced supervised GAN training.
+More recently, RareGAN (Lin et al., 2022) addresses unlabeled long-tailed data through a weighted loss and adaptive labeling budgets, improving both balance and generative diversity.
+Despite these advances, GANs remain difficult to train reliably, especially under strong imbalance.
+Diffusion models have emerged as a more stable alternative to GANs, achieving superior image quality and convergence behavior (Dhariwal & Nichol, 2021).
+Several diffusion-based methods explicitly tackle data imbalance.
+Class-Balancing Diffusion Models (Qin et al., 2023) modify the denoising process to be class-invariant, while (Zhang et al., 2024) employs weighted score matching with Bayesian calibration to transfer knowledge from majority to minority classes, improving diversity and fidelity on long-tailed datasets.
+Heavy-Tailed Diffusion Models (Pandey et al., 2025) take a different approach by replacing the Gaussian noise assumption with a Student’s $t$ formulation, offering a more robust fit for imbalanced distributions.
+Nevertheless, effectively addressing class imbalance within diffusion frameworks remains an open and active research area.
+Variational Autoencoders (VAEs) (Kingma & Welling, 2013) are valuable models when interpretable, structured latent spaces and efficient single-step sampling are needed.
+VAEs provide a principled probabilistic foundation, stable training dynamics, and compatibility with latent-variable pipelines such as latent diffusion (Rombach et al., 2022).
+However, in class imbalanced setting, standard VAEs using isotropic Gaussian priors struggle to capture heavy-tailed or rare structures, creating class-coverage bottlenecks (Tam & Dunson, 2025).
+Therefore, prior work explores non-Gaussian priors, especially Student’s $t$-distributions (Takahashi et al., 2018; Abiri & Ohlsson, 2020; Eguchi, 2021; Kim et al., 2024), to improve robustness to outliers and class imbalance.
+In this work, we view class imbalance not only as a sampling issue but as a geometric one: optimizing under the empirical distribution tends to allocate latent probability mass in proportion to class frequency.
+Consequently, majority classes occupy larger effective regions of latent space, while minority classes are confined to smaller regions with reduced representational capacity.
+This interpretation is particularly transparent in latent-variable models with an explicit global prior, such as VAEs, where representational allocation is directly governed by prior density.
+While imbalance also affects GANs and diffusion models, in those frameworks it is intertwined with adversarial training dynamics or iterative score-based denoising.
+In contrast, VAEs allow us to study how prior specification alone shapes latent geometry.
+We therefore focus on VAEs to isolate the role of prior-induced frequency bias in a controlled and analytically tractable setting.
+Current VAE based approaches retain a single global prior (Kingma & Welling, 2013; Takahashi et al., 2018; Mathieu et al., 2019b; Abiri & Ohlsson, 2020).
+Hence through our geometric view the latent remains frequency-aligned where class regions scale with empirical probability and heavy tails alone do not eliminate this frequency-induced geometric bias.
+Indeed, optimizing a VAE under $p_{data}(x)=\sum_{y}p(y)p(x|y)$ tends to allocate latent regions whose effective volume correlates with class frequency $p(y)$.
+We address this issue by introducing the Conditional-$t^{3}$VAE (C-$t^{3}$VAE), which imposes a per-class Student’s $t$-distribution prior over the joint latent-output space.
+The design promotes uniform prior mass across class-conditioned components under the sampling distribution, mitigating majority-class dominance while the heavy tails capture intra-class variation.
+For class-balanced generation, we derive an equal-weight latent mixture of Student’s $t$-distributions with analytically derived component variances.
+Importantly, C-$t^{3}$VAE class-specific heavy-tailed priors induce a structured latent mixture whose geometry differs fundamentally from both global heavy-tailed priors and Gaussian conditional VAEs.
+We summarize our main contributions in the following points :
   • We propose the C-$t^{3}$VAE model with a training objective based on the $\gamma$-power divergence.
   • We develop an equal-weight latent mixture sampling scheme with analytically derived optimal variance scaling for each component.
-  • We outperform relevant baselines in FID on SVHN-LT [11], CIFAR100-LT [12], and CelebA [13] under severe imbalance, and show via per-class evaluation that C-$t^{3}$VAE better avoids mode collapse, exceeding a conditional VAE in per class Recall and F1 while remaining competitive on Precision.
-  • We identify the imbalance ratio threshold $\rho\approx 3$, beyond which Gaussian priors become suboptimal, providing guidance for model selection on skewed datasets.
-2 Related Work
-Since the introduction of VAEs [4], many extensions have sought to improve latent representation by replacing the standard Gaussian prior with more expressive alternatives.
-These include Gaussian mixtures [14, 15], hyperspherical priors [16], normalizing flows [17], Riemannian priors [18], and implicit distributions [19].
-While most retain the Evidence Lower Bound (ELBO) formulation, others adopt alternative objectives or divergence measures for added flexibility.
-To address long-tailed or imbalanced data, Student’s t-distributions have been explored for their robustness and heavy tails [6].
-Methods such as [7] and [8] model the latent space of the autoencoder through a t-distributed prior, and rely on KL-divergence-based ELBO objectives.
-However, since the KL divergence lacks a closed-form solution for the Student’s t-distributions, these methods resort to numerical approximation.
-The $t^{3}$VAE [10] improves on this by modeling the joint latent-output distribution and optimizing a closed-form $\gamma$-divergence objective [9].
-Nonetheless, it still employs a global latent prior, which results in a latent space volume allocation reflecting class frequency and leading to imbalance present in the generated samples.
-Other works address fairness in generative modeling using normalizing flows [17] and diffusion models [20, 21].
-However, since leading image generation models are based on Latent Diffusion Models [5] and VAEs are a corner-stone in the design of these models, we deem it crucial to improve VAEs’ ability to handle imbalanced data.
-Consequently, as fairness in latent space allocation remains under-explored in VAEs, we introduce a class-conditional, heavy-tailed prior to address this gap and allow for balanced latent space allocation across classes.
+  • We outperform relevant baselines in FID on SVHN-LT (Netzer et al., 2011), CIFAR100-LT (Cao et al., 2019), and CelebA (Liu et al., 2015) under severe imbalance, and show via per-class evaluation that C-$t^{3}$VAE better avoids mode collapse, exceeding a Gaussian conditional VAE in per class Recall and F1 while remaining competitive on Precision.
+  • We empirically observe a transition regime around $\rho\approx 5$, beyond which Gaussian priors become suboptimal, providing guidance for model selection on skewed datasets.
+2 Related works
+Since the introduction of Variational Autoencoders (VAEs) (Kingma & Welling, 2013), numerous extensions have aimed to enhance latent representation expressiveness by replacing the standard Gaussian prior with more flexible formulations.
+Notable examples include Gaussian mixtures (Dilokthanakul et al., 2016; Saseendran et al., 2021), hyperspherical priors (Davidson et al., 2018), normalizing flows (Jaini et al., 2020), Riemannian manifolds (Chadebec et al., 2023), and implicit distributions (Takahashi et al., 2019).
+Most of these retain the Evidence Lower Bound (ELBO) optimization framework, while others introduce alternative divergence measures for greater modeling flexibility (Makhzani et al., 2016).
+Hierarchical VAEs (Vahdat & Kautz, 2020) and vector-quantized VAEs (van den Oord et al., 2017) have also been proposed to improve disentanglement and mitigate posterior collapse.
+Student’s $t$-distributions have been widely studied for their heavy-tailed robustness to outliers (Tam & Dunson, 2025).
+Early works (Mathieu et al., 2019a; Abiri & Ohlsson, 2020) incorporated $t$-distributed priors into the VAE framework using KL-divergence-based ELBO objectives to promote robust latent encodings.
+However, because the KL divergence between $t$-distributions lacks a closed-form solution, such approaches depend on numerical approximations, increasing computational cost.
+(Takahashi et al., 2018) proposes a Student’s $t$ decoder to improve generative performance, but the heavy-tailed assumption is confined to the decoder, leaving the latent space Gaussian and limiting validation to tabular data.
+The $t^{3}$VAE (Kim et al., 2024) further advances this line of work by jointly modeling latent and output distributions under Student’s $t$-assumptions for the encoder, decoder, and prior, replacing the KL divergence with a closed-form $\gamma$-divergence (Eguchi, 2021).
+While several works improve prior expressiveness or robustness, they do not explicitly analyze how global prior formulations induce frequency-aligned allocation of latent probability mass under imbalanced data.
+In particular, existing approaches focus on better density modeling or optimization stability, rather than controlling how representational capacity is distributed across classes.
+As a result, even expressive or heavy-tailed priors may continue to reflect empirical class proportions in latent space.
+A complementary line of research addresses imbalance through conditional modeling.
+Conditional VAEs (CVAEs) (Kingma et al., 2014; Sohn et al., 2015) condition latent encodings on class labels, enabling targeted generation of minority categories.
+However, their Gaussian priors poorly approximate heavy-tailed or rare data structures.
+Our C-$t^{3}$VAE bridges two threads: it inherits the heavy-tailed robustness of $t^{3}$VAE while adopting the class-conditional structure of CVAEs, and further introduces a theoretically grounded equal-weight mixture sampling scheme that explicitly counteracts majority-class dominance at generation time.
+Hence, our contribution is not merely heavy tails nor conditioning, but decoupling representational capacity from empirical frequency through prior mass allocation.
 3 Background
 This section introduces the theoretical background and baseline models relevant to our work.
-We assume access to a labeled, imbalanced dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}\in\mathbb{R}^{n}$ is a data sample of dimension $n$, $y_{i}\in\{1,\dots,K\}$ its class label and $m$ the latent space dimension.
-3.1 VAEs and Conditional-VAEs
-VAEs [4] are generative models trained via variational inference by maximizing the ELBO of the log-likelihood. The standard objective of this model is
+We assume access to a labeled, imbalanced dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}\in\mathbb{R}^{n}$ is a data sample of dimension $n$, $y_{i}\in\{1,\dots,K\}$ its class label with $K$ being the number of classes and $m$ being the latent space dimension.
+3.1 VAEs and Conditional VAEs
+VAEs (Kingma & Welling, 2013) are generative models trained via variational inference by maximizing the Evidence Lower Bound (ELBO) of the log-likelihood. The standard objective of this model is
 $\mathcal{L}_{\theta,\phi}:=\mathbb{E}_{z\sim q_{\phi}(\cdot|x)}[\log p_{\theta}(x|z)]-\mathcal{D}_{KL}(q_{\phi}(z|x)\|p(z)),$ (1)
-where the first term is the reconstruction loss with $p_{\theta}(x|z)$ being the decoder.
+where the first term is the reconstruction loss with $p_{\theta}(x|z)$ being the decoder model.
 The second term is the Kullback–Leibler (KL) divergence between the approximate posterior $q_{\phi}(z|x)$ and the prior $p(z)$.
-The $\beta$-VAE is a weighted variant of the VAE model which introduces a $\beta$ scaling term for the KL divergence [22]:
+The $\beta$-VAE is a weighted variant of the VAE model which introduces a $\beta$ scaling term for the KL divergence (Higgins et al., 2017):
 $\mathcal{L}_{\theta,\phi}:=\mathbb{E}_{z\sim q_{\phi}(\cdot|x)}[\log p_{\theta}(x|z)]-\beta\mathcal{D}_{KL}(q_{\phi}(z|x)\|p(z)),$ (2)
 with $p(z)\sim\mathcal{N}_{m}(0,I)$, $q_{\phi}(\cdot|x)\sim\mathcal{N}_{m}(\mu_{\phi}(x),\Sigma_{\phi}(x))$, and $p_{\theta}(x|z)\sim\mathcal{N}_{m}(\mu_{\theta}(z),\sigma^{2}I)$.
-$\mu_{\phi}(\cdot)$ and $\Sigma_{\phi}(\cdot)$ are the mean and the covariance matrices inferred through a neural network with parameter $\phi$ given the input $x$.
+$\mu_{\phi}(\cdot)$ and $\Sigma_{\phi}(\cdot)$ are the mean and covariance matrices of the approximate posterior.
+They are inferred by a neural network with parameters $\phi$ given the input $x$.
+Moreover, $\mu_{\theta}(\cdot)$ is the decoder neural network with parameter $\theta$ and $\sigma$ is a parameter controlling the decoder’s output covariance.
 This variant of the VAE model allows to place more weight on disentangling the latent space or on the reconstruction of the data points.
@@
 (1) and Eq.
-(2) optimize the ELBO over the data distribution $p_{data}(x)$, which can be decomposed as $p_{\mathrm{data}}(x)=\sum_{y_{i}}p(y_{i})\,p_{\mathrm{data}}(x\mid y_{i}),$ this optimization inherently biases the model toward head classes with larger $p(y_{i})$.
+(2) optimize the ELBO over the data distribution $p_{data}(x)$, which can be decomposed as $p_{\mathrm{data}}(x)=\sum_{y_{i}}p(y_{i})\,p_{\mathrm{data}}(x\mid y_{i})$.
+In the context of imbalanced data, this optimization inherently biases the model toward head classes with larger $p(y_{i})$.
 As a result, most generated samples come from overrepresented classes, while tail classes’ samples are underrepresented and of lower quality, a phenomenon commonly referred to as mode collapse.
-Therefore, when labels are available, it is preferable to define class-conditional approximate posterior and prior distributions: $q_{\phi}(z|x,y)$ and $p(z|y)$.
-This yields the Conditional-VAE (CVAE) model trained using the objective [23]:
+Therefore, when labels are available, it is preferable to define class-conditional posterior and prior distributions: $q_{\phi}(z|x,y)$ and $p(z|y)$.
+This yields the Conditional-VAE (CVAE) model trained using the objective (Kingma et al., 2014):
 $\sum_{y}\ \mathbb{E}_{z\sim q_{\phi}(\cdot|x,y)}[\log p_{\theta}(x|z,y)]-\beta\mathcal{D}_{KL}(q_{\phi}(z|x,y)\|p(z|y)).$ (3)
-Here, in Eq.
-(3) we constrain all $p(y_{i})$ to be equal and omit them from the loss function in order not to exacerbate the issue of class imbalance in the latent space.
+In practice, in Eq.
+(3) we optimize the class-conditional objective with uniform weighting across labels, removing explicit frequency-dependent scaling from the loss.
 Also, we define $p(z|y)\sim\mathcal{N}_{m}(\mu_{y},I)$ with learnable class-wise means $\mu_{y}$.
@@
 Nevertheless, despite conditioning, this formulation remains Gaussian.
-Unlike Student’s t-distributions, Gaussian priors poorly approximate heavy-tailed data distributions [6].
-3.2 Multivariate Student’s t-Distribution and $\gamma$-Power Divergence
+Unlike Student’s $t$-distributions, Gaussian priors poorly approximate heavy-tailed data distributions (Tam & Dunson, 2025).
+3.2 Multivariate Student’s $t$-Distribution
 A $d$-dimensional Student’s $t$-distribution with mean $\mu\in\mathbb{R}^{d}$, covariance $\Sigma\in\mathbb{R}^{d\times d}$, and degrees of freedom $\nu>2$ is a heavy-tail, super-Gaussian distribution defined as
@@
 The power form of this distribution prevents a closed-form KL divergence between two Student’s $t$-distributions.
-Instead, the $\gamma$-power divergence $\mathcal{D}_{\gamma}(q\|p)$ is used [9, 10].
-This divergence is defined for $q\sim t_{d}(\mu_{0};\Sigma_{0};\nu)$, $p\sim t_{d}(\mu_{1};\Sigma_{1};\nu)$ starting from the $\gamma$-entropy $\mathcal{H}_{\gamma}(p)$ and cross-entropy $\mathcal{C}_{\gamma}(q,p)$
+Instead, the $\gamma$-power divergence $\mathcal{D}_{\gamma}(q\|p)$ is used (Eguchi, 2021; Kim et al., 2024).
+This divergence is defined for $q\sim t_{d}(\mu_{0};\Sigma_{0};\nu)$, $p\sim t_{d}(\mu_{1};\Sigma_{1};\nu)$ as
+$\mathcal{D}_{\gamma}(q\|p):=\frac{\mathcal{C}_{\gamma}(q,p)-\mathcal{H}_{\gamma}(p)}{\gamma}$ (5)
+with $\gamma=-\frac{2}{\nu+d}$, the $\gamma$-entropy $\mathcal{H}_{\gamma}(p)$ and $\gamma$-cross-entropy $\mathcal{C}_{\gamma}(q,p)$ being
 $\mathcal{H}_{\gamma}(p):=-\|p\|_{1+\gamma}=-\left(\int p(x)^{1+\gamma}dx\right)^{\frac{1}{1+\gamma}},\qquad\mathcal{C}_{\gamma}(q,p):=-\int q(x)\left(\frac{p(x)}{\|p\|_{1+\gamma}}\right)^{\gamma}dx.$
-$\mathcal{D}_{\gamma}(q\|p):=\gamma^{-1}\left(\mathcal{C}_{\gamma}(q,p)-\mathcal{H}_{\gamma}(p)\right)$ (5)
-with $\gamma=-\frac{2}{\nu+d}$.
 Then, substituting the definition of a Student’s $t$-distribution from Eq.
 (4) into Eq.
-(5), the following closed-form formula for the $\gamma$-power divergence can be derived (See Appendix A):
+(5), the following closed-form formula for the $\gamma$-power divergence can be derived (Derivation of Proposition 3 in (Kim et al., 2024))
 $\mathcal{D}_{\gamma}(q\|p)$ $=-\frac{C_{\nu,d}^{\frac{\gamma}{1+\gamma}}}{\gamma}\left(1+\frac{d}{\nu-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg[-|\Sigma_{0}|^{-\frac{\gamma}{2(1+\gamma)}}\quad\left(1+\frac{d}{\nu-2}\right)$
-$\quad+|\Sigma_{1}|^{-\frac{\gamma}{2}}|\Sigma_{0}|^{\frac{\gamma^{2}}{2(1+\gamma)}}\Bigg{(}1+\frac{\operatorname{Tr}\left(\Sigma_{1}^{-1}\Sigma_{0}\right)}{\nu-2}+\frac{(\mu_{0}-\mu_{1})^{\top}\Sigma_{1}^{-1}(\mu_{0}-\mu_{1})}{\nu}\Bigg{)}\Bigg{]}.$ (6)
+$\quad+|\Sigma_{1}|^{-\frac{\gamma}{2(1+\gamma)}}\Bigg(1+\frac{\operatorname{Tr}\left(\Sigma_{1}^{-1}\Sigma_{0}\right)}{\nu-2}+\frac{(\mu_{0}-\mu_{1})^{\top}\Sigma_{1}^{-1}(\mu_{0}-\mu_{1})}{\nu}\Bigg)\Bigg].$ (6)
 3.3 $t^{3}$-Variational Autoencoder
 3.3.1 Definition
-The $t^{3}$VAE model [10] is a non-ELBO-based autoencoder which models the joint prior distribution $p_{\theta}(x,z)$ using multivariate Student’s t-distributions
-$p_{\theta}(x,z)\propto\sigma^{-n}\left[1+\frac{1}{\nu}\left(\|z\|^{2}+\frac{\|x-\mu_{\theta}(z)\|^{2}}{\sigma^{2}}\right)\right]^{-\frac{\nu+m+n}{2}},$
-where $\mu_{\theta}(\cdot)$ is the decoder neural network with parameter $\theta$ whereas $\sigma$ is a parameter controlling the decoder’s output covariance.
-From this joint distribution, the marginal latent prior $p(z)$ and decoder distribution $p_{\theta}(x|z)$ can be defined.
-Furthermore, the approximate posterior distribution is defined as :
+The $t^{3}$VAE model (Kim et al., 2024) is a non-ELBO-based autoencoder which models the joint prior distribution $p_{\theta}(x,z)$ using multivariate Student’s $t$-distribution
+$p_{\theta}(x,z)=\sigma^{-n}C_{\nu,m+n}\left[1+\frac{1}{\nu}\left(\|z\|^{2}+\frac{\|x-\mu_{\theta}(z)\|^{2}}{\sigma^{2}}\right)\right]^{-\frac{\nu+m+n}{2}}.$
+From this joint distribution, the marginal latent prior $p(z)$ and decoder distribution $p_{\theta}(x|z)$ can be defined as follows
+$p(z)=t_{m}(z|0,I,\nu),\quad p_{\theta}(x|z)=t_{n}\left(x\middle|\mu_{\theta}(z),\frac{1+\nu^{-1}\|z\|^{2}}{1+\nu^{-1}m}\sigma^{2}I,\nu+m\right).$
+Furthermore, the posterior distribution is defined as :
 $q_{\phi}(z|x)=t_{m}\left(x\middle|\mu_{\phi}(x),\frac{\Sigma_{\phi}(x)}{1+\nu^{-1}n},\nu+n\right).$
@@
 $\mathcal{L}_{\gamma}=\mathbb{E}_{x}\Bigg[\frac{\mathbb{E}_{z}\left[\|x-\mu_{\theta}(z)\|^{2}\right]}{\sigma^{2}}+\|\mu_{\phi}(x)\|^{2}+\frac{\nu\operatorname{Tr}\left(\Sigma_{\phi}(x)\right)}{\nu+n-2}-\frac{\nu C_{1}}{C_{2}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}\Bigg],$ (7)
-with $\gamma=-\frac{2}{\nu+n+m}$ and $C_{1}$ and $C_{2}$ being theoretically derived constants.
+with $\gamma=-\frac{2}{\nu+n+m}$ and $C_{1}$ and $C_{2}$ being constants theoretically derived in (Kim et al., 2024).
 We note that the first term in this loss function represents the standard reconstruction term in VAE models and the rest of the terms are regularization terms over the latent space.
-To sample from the latent space of the $t^{3}$VAE, [10] propose the $p_{\nu}^{\star}(z)=t_{m}(0,\tau^{2}I,\nu+n)$ distribution with
-$\tau^{2}=\frac{1}{1+\nu^{-1}n}\left(\frac{C_{\nu,n}}{\sigma^{n}}\cdot\frac{\nu-2}{\nu+n-2}\right)^{\frac{2}{\nu+n-2}}.$ (8)
+To sample from the latent space of the $t^{3}$VAE, (Kim et al., 2024) propose the $p_{\nu}^{\star}(z)=t_{m}(0,\tau^{2}I,\nu+n)$ distribution with
+$\tau^{2}=\left(1+\nu^{-1}n\right)^{-1}\left(\sigma^{n}C^{-1}_{\nu,n}\quad\frac{\nu+n-2}{\nu-2}\right)^{-\frac{2}{\nu+n-2}}.$ (8)
 Moreover, sampling from a multi-variate Student’s $t$-distribution $T\sim t_{d}(\mu,\Sigma,\nu)$ both in the learning (Eq.
@@
 (8)) phases is performed through the standard reparameteration trick for Student’s $t$-distributions $T:=\mu+Z\sqrt{\nu V^{-1}}$ where $Z\sim\mathcal{N}(0,\Sigma)$ and $V\sim\mathcal{X}^{2}(\nu)$.
-3.3.2 $\beta$-$t^{3}$VAE
+3.3.2 $\beta$-$t^{3}$VAE extension
 From Eq.
@@
 Similarly to $\beta$-VAE models, this improves the versatility of the model and allows either a focus on generation or disentangling.
-3.3.3 $\tau^{2}$ Improvement
-Closely analyzing the proposed derivation of $\tau^{2}$ of the $t^{3}$VAE model, we discovered a subtle issue in its mathematical formulation.
-The employed $\gamma$-power divergence presents a discrepancy to the correct formula in Eq.
-(6).
-We revised the formulation and the corrected $\tau^{2}$ is (See Appendix D):
-$\tau^{2}$ $=\frac{1}{1+\nu^{-1}n}\left(|\Sigma_{\phi}(x)|^{\frac{\gamma}{2}}\frac{C_{\nu,n}}{\sigma^{n}}\cdot\frac{\nu-2}{\nu+n-2}\right)^{\frac{2\gamma}{(1+\gamma)(2+\gamma m)}}\approx\frac{1}{1+\nu^{-1}n}\left(\frac{C_{\nu,n}}{\sigma^{n}}\cdot\frac{\nu-2}{\nu+n-2}\right)^{-\frac{2\gamma}{2+\gamma m}}.$ (9)
-The corrected exact form of $\tau^{2}$ is applicable when $|\Sigma_{\phi}(x)|$ is known and when the dimension of the data is low.
-However, for high dimensional data, as handled in this work, we get $\gamma\approx 0$.
-Hence, one can use the approximation without any loss in accuracy.
-We note that the new form of $\tau^{2}$ leads to a similar empirical value of standard deviation compared to the previous form.
-Nevertheless, for correctness, in our sampling from the latent space of the $t^{3}$VAE model we use the approximation in Eq.
-(9).
-In summary, although the $t^{3}$VAE effectively models heavy-tailed distributions through Student’s t-distributions and $\gamma$-power divergence, it does not explicitly address class imbalance in the latent space by not allocating equal volume for each class.
-In the next section, we introduce a class-conditional variant of the $t^{3}$VAE, designed to ensure fair and balanced generation across all classes.
+In summary, although the $t^{3}$VAE effectively models heavy-tailed distributions through Student’s $t$-distributions and $\gamma$-power divergence, it does not explicitly address class imbalance in the latent space as it does not allocate equal prior mass across class-conditioned components.
+In the next section, we introduce a class-conditional variant of the $t^{3}$VAE, designed to enable class-balanced generation across all classes.
 4 Conditional $t^{3}$-Variational Autoencoder
-We propose the Conditional $t^{3}$-Variational Autoencoder (C-$t^{3}$VAE), present its formulation, training objective, and sampling strategy.
-C-$t^{3}$VAE models the latent space as a mixture of Student’s t-distributions, one per class, ensuring equal latent volume allocation and promoting fairness in generation.
-Intra-class imbalance is further handled by the heavy-tailed nature of the Student’s t-distribution prior.
+In this section, we propose the Conditional $t^{3}$-Variational Autoencoder (C-$t^{3}$VAE), present its formulation, training objective, and sampling strategy.
+C-$t^{3}$VAE models the latent space as a mixture of Student’s $t$-distributions, one per class, inducing uniform prior mass across class-conditioned components under the sampling distribution.
+Intra-class variability is captured through the heavy-tailed nature of the Student’s $t$ prior.
 4.1 Model definition
@@
 Moreover, $\Sigma_{x}$ and $\Sigma_{y}$ are the covariance matrices of the prior distributions over the latent and output variables.
-From this joint distribution, we can derive the conditional latent prior $p(z|y)=t_{m}(z|\mu_{y},\Sigma_{y},\nu)$ and decoder distribution (See Appendix B)
-$p_{\theta}(x|z,y)=t_{n}\Bigg{(}x\Bigg{|}\mu_{\theta}(z),\frac{\left(1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right)}{(1+\nu^{-1}m)}\Sigma_{x},\nu+m\Bigg{)}.$
-Furthermore, as in $t^{3}$VAE, we define the approximate posterior $q_{\phi}(z|x)$ as a multivariate Student’s t-distribution capturing heavy-tailed structure in the latent space :
+From this joint distribution, we can derive the conditional latent prior $p(z|y)$ and decoder distribution $p_{\theta}(x|z,y)$ (See Appendix A)
+$p(z|y)=t_{m}(z|\mu_{y},\Sigma_{y},\nu),\quad p_{\theta}(x|z,y)=t_{n}\Bigg(x\Bigg|\mu_{\theta}(z),\frac{\left(1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right)}{(1+\nu^{-1}m)}\Sigma_{x},\nu+m\Bigg).$
+By defining class-specific priors, the model ensures that each class occupies a proportionally equal latent region, countering the frequency-aligned bias of a global prior.
+Furthermore, as in $t^{3}$VAE, we define the posterior $q_{\phi}(z|x)$ as a multivariate Student’s $t$-distribution capturing heavy-tailed structure in the latent space :
 $q_{\phi}(z|x)=t_{m}\left(z\middle|\mu_{\phi}(x),\frac{\Sigma_{\phi}(x)}{1+\nu^{-1}n},\nu+n\right).$
+Although the posterior is defined as $q_{\phi}(z|x)$ without explicit class conditioning, the class-specific prior and objective derived in the following section enforce class-discriminative latent encodings.
 4.2 Objective function
-Harnessing Eq. (5) and the defined prior and posterior distributions of the proposed C-$t^{3}$VAE, we derive in Appendix C the following class-wise objective
+Harnessing Eq. (5) and the defined prior and posterior distributions of the proposed C-$t^{3}$VAE, we derive in Appendix B the following class-wise objective
 $\mathcal{L}(\gamma,y)=\mathbb{E}_{x}$ $\Bigg[\mathbb{E}_{z}\Big[(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))\Big]+(\mu_{\phi}(x)-\mu_{y})^{\top}\Sigma_{y}^{-1}(\mu_{\phi}(x)-\mu_{y})$
@@
 $\mathcal{L}(\gamma,y)=\mathbb{E}_{x}\Bigg[\frac{\mathbb{E}_{z}\left[\|x-\mu_{\theta}(z)\|^{2}\right]}{\sigma^{2}}+\|\mu_{\phi}(x)-\mu_{y}\|^{2}+\frac{\nu\operatorname{Tr}\left(\Sigma_{\phi}(x)\right)}{\nu+n-2}-\frac{\nu C_{1}}{C_{2}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}\Bigg].$ (9)
-Therefore, we express the final loss function $\mathcal{L}(\gamma)$ over the whole dataset as : $\mathcal{L}(\gamma)=\sum_{y}\mathcal{L}(\gamma,y).$ As in Eq.
-(3), here too we consider all $p(y_{i})$ to be equal and hence rule out their contribution to the loss function.
-This is done to avoid emphasizing the imbalance present in the data.
+From Eq.
+(9), we observe that the first term aims to minimize the input reconstruction error, ensuring data fidelity, while the second term promotes alignment between the input class mean and the inferred posterior mean.
+The third and fourth terms jointly regularize the posterior covariance $\Sigma_{\phi}(x)$; specifically, the trace penalty restricts the total variance to encourage compactness, whereas the determinant term acts as a barrier function to prevent the covariance matrix from collapsing to a singularity.
+Finally, we express the final loss function $\mathcal{L}(\gamma)$ over the whole dataset as :
+$\mathcal{L}(\gamma)=\sum_{y}\mathcal{L}(\gamma,y).$
+As in Eq. (3), here too we use uniform class weighting so the loss does not scale with empirical frequency and the generative step targets all classes uniformly.
 4.3 Sampling distribution
-Similarly to the objective function of $t^{3}$VAE, $\mathcal{L}(\gamma,y)$ in Eq.
-(10) can be decomposed into a reconstruction term and regularization terms.
-To sample from the latent space of the C-$t^{3}$VAE, we focus on the regularization terms and define the following sampling distribution :
-$p_{\nu}^{\star}(z)=\sum_{y=1}^{K}\alpha_{y}\cdot t_{m}(\mu_{y},\tau^{2}I,\nu+n),\quad\forall y,\;\;\alpha_{y}=\frac{1}{K}.$ (11)
-The theoretical derivation of the variance $\tau^{2}$ leads to the form expressed in Eq. (9) (Derivation in Appendix D).
+To sample from the latent space of the C-$t^{3}$VAE, we define the following sampling distribution :
+$p_{\nu}^{\star}(z)=\sum_{y=1}^{K}\alpha_{y}\cdot p_{\nu,y}^{\star}(z)=\sum_{y=1}^{K}\alpha_{y}\cdot t_{m}(\mu_{y},\tau^{2}I,\nu+n),\quad\forall y,\;\;\alpha_{y}=\frac{1}{K}.$ (10)
+For the the variance parameter $\tau^{2}$ in $t_{m}(\mu_{y},\tau^{2}I,\nu+n)$, we derive it theoretically.
+In this derivation we first derive the $\gamma$-power divergence between $p_{\nu,y}^{\star}(z)=t_{m}(\mu_{y},\tau^{2}I,\nu+n)$ and $q_{\phi}(z|x)$ for every $y$.
+Then, we compare the obtained result to the corresponding regularization terms in $\mathcal{L}(\gamma,y)$ Eq.
+(9) (See Appendix C for details).
+The obtained form is similar to the form expressed in Eq.
+(8).
 The mixture-based sampling distribution we define in Eq.
-(11) with equal $\alpha_{y}$ ensures a uniformly sampled synthetic data across all classes, regardless of their frequency in the original training data.
-As a result, C-$t^{3}$VAE equipped with this sampling distribution mitigates the common problem in generative models where head-class samples dominate due to their density in latent space.
-Furthermore, by modifying the mixture weights $\alpha_{y}$, one can prioritize specific classes.
-This makes our method flexible for targeted data augmentation or fairness-aware sampling strategies.
-4.3.1 $\beta$-C-$t^{3}$VAE
+(10) with equal $\alpha_{y}$ is not a post-hoc balancing heuristic.
+Rather, it is the generative counterpart of promoting uniform prior mass across class-conditioned latent components.
+Sampling proportionally to empirical frequency would partially reintroduce frequency-aligned bias at generation time, attenuating the structural change imposed during training.
+Our framework also allows prioritization by modifying the mixture weights $\alpha_{y}$ when targeted generation is desired; however, in this work we focus on the balanced setting.
+4.3.1 $\beta$-C-$t^{3}$VAE extension
 As with $t^{3}$VAE, the class-wise objective defined in Eq.
@@
 By preceding the regularization term with a $\beta$ scalar, we can define a $\beta$-C-$t^{3}$VAE model thereby improving the domain of applicability of the model.
-Overall, C-$t^{3}$VAE provides a principled, flexible, and tractable framework for fair generative modeling, particularly under class-imbalanced conditions.
-In the following, we study the C-$t^{3}$VAE’s performance across multiple datasets with varying imbalance degrees.
 5 Experiments
-In this section we outline the generative performance of the proposed C-$t^{3}$VAE model on labeled datasets compared to relevant VAE baselines [^1] We restrict our comparisons to VAE-based generative models, as Latent Diffusion Models [5], the state-of-the-art in image generation, depend on VAEs as a core component.
-Enhancing fairness in VAEs under class imbalance is therefore a necessary precursor to advancing more complex models..
-We conduct experiments on three datasets notably SVHN-LT [11], CIFAR100-LT [24, 12] and CelebA [13] each chosen to highlight different challenges.
-5.1 Evaluation procedure
-All models are evaluated using Fréchet Inception Distance (FID) [25], computed against a balanced test set for each dataset.
-This setup measures how effectively a model overcomes training set imbalance by assessing its ability to generate high-quality samples across all classes.
-To evaluate robustness, we impose varying degrees of imbalance during training.
-For SVHN-LT and CIFAR100-LT, we introduce class imbalance by applying an exponential decay to the number of samples per class after equalizing class sizes in the original train and test sets.
-The imbalance ratio $\rho$ defines the ratio between the most and least frequent classes, with class-wise sample counts $M_{y_{i}}$ given by: $M_{y_{i}}=M\cdot\rho^{-\frac{y_{i}-1}{K-1}},$ where $M$ is the original per-class sample count.
-For CelebA, we compute FID per attribute, treating each attribute (eg.
-Mustache) and its negation (eg.
-no Mustache) as separate binary classes.
-The training set uses CelebA’s inherent imbalance, while the test set is balanced by downsampling to the smaller class size.
-Multi-attribute generation is not considered in this work and is left for future exploration.
-5.2 Results
-We present both quantitative and qualitative results of the C-$t^{3}$VAE model and the models it directly improves upon notably the VAE, C-VAE and $t^{3}$VAE models with their $\beta$ variants.
-This controlled comparison helps isolate the contributions of key design choices:
+This section presents quantitative and qualitative results for the C-$t^{3}$VAE model and closely related baselines (VAE, C-VAE, and $t^{3}$VAE, with their $\beta$ variants).
+This controlled comparison isolates the contributions of our design choices:
   • VAE : ELBO trained standard Gaussian-based VAEs.
   • C-VAE : VAE supplemented by conditional Gaussian priors to assess the class conditioning effect without changing the prior family.
-  • $t^{3}$VAE : Student’s t-distribution latent prior and $\gamma$-power divergence objective, does not use class-conditional priors, nor allows for class conditional generation.
-Through a comparison with this model we assess the role of conditional modeling on Student’s t-distribution priors.
-In the following, we analyze the latent sampling standard deviation $\tau$ employed in $t^{3}$VAE and C-$t^{3}$VAE models, varying it to assess alignment between empirical and theoretical values.
-Then, we present FID comparisons across baselines with optimized hyper-parameters (The tuning of $\beta$, $\nu$ and $\tau$ hyper-parameters is reported in Appendix F), followed by per-class generative evaluation.
+  • $t^{3}$VAE : Student’s $t$-distribution latent prior and $\gamma$-power divergence objective; it does not use class-conditional priors and does not allow class-conditional generation.
+Comparing to this model isolates the effect of conditional modeling with heavy-tailed priors.
+We also evaluate $\beta$-VAE, $\beta$-C-VAE, and $\beta$-$t^{3}$VAE variants to assess the effect of tuning regularization on latent disentanglement and generative performance.
+Besides, we note that we do not aim to compete with state-of-the-art diffusion or adversarial models in raw sample quality.
+Instead, we isolate the effect of latent prior geometry under class imbalance within a controlled variational framework.
+Comparing across fundamentally different architectural families would confound inductive bias, training stability, and objectives with the geometric contribution we study.
+Hence, our evaluation is restricted to VAE-family models and we do not benchmark against diffusion-based generators.
+In the following, we first present the evaluation metrics and protocol, then analyze the latent sampling standard deviation $\tau$ used in $t^{3}$VAE and C-$t^{3}$VAE by varying $\tau$ to assess alignment between empirical and theoretical values.
+We then report FID comparisons across baselines with optimized hyper-parameters (the tuning of $\beta$, $\nu$, and $\tau$ is reported in Appendix E), followed by per-class generative evaluation.
+5.1 Evaluation metrics
+In this section, we present the evaluation metrics used throughout this work.
+We report FID and generative Precision, Recall and F1 because they are the most commonly used metrics in the VAE and long-tailed generation literature, enabling direct comparison with prior work.
+Also, these metrics allow us to assess whether the class-conditional heavy-tailed prior in C-$t^{3}$VAE improves minority-class generation relative to baselines.
+5.1.1 Fréchet Inception Distance
+The Fréchet Inception Distance (FID) (Heusel et al., 2017) is a standard metric for evaluating the quality of synthetic images.
+It measures the similarity between the distributions of real and generated images, with a score of 0 indicating identical distributions.
+To compute the FID, feature encodings of real and generated images are extracted using the InceptionV3 network (Szegedy et al., 2016), excluding the classification layer.
+Assuming these encodings follow a multivariate Gaussian distribution, the mean vectors $\mu_{r}$ and $\mu_{s}$ and covariance matrices $\Sigma_{r}$ and $\Sigma_{s}$ are estimated for the real and synthetic image sets, respectively.
+The FID is then calculated as:
+$FID=\|\mu_{r}-\mu_{s}\|_{2}^{2}+\operatorname{Tr}\left(\Sigma_{r}+\Sigma_{s}-2\cdot\left(\Sigma_{r}\Sigma_{s}\right)^{\frac{1}{2}}\right).$
+5.1.2 Generative Precision, Recall, and F1
+While FID provides a holistic measure of distributional similarity, it does not explicitly assess the quality and diversity of generated samples.
+To address this, generative Precision and Recall metrics have been proposed (Kynkäänniemi et al., 2019).
+These metrics evaluate respectively the sharpness and mode coverage of generated samples relative to the target distribution.
+Their harmonic mean, the F1 score, offers a balanced assessment.
+Given sets of real and generated samples $X_{r}$ and $X_{g}$, feature representations are extracted using a classifier network, yielding vectors $\delta_{r}$ and $\delta_{g}$.
+The complete sets of feature vectors are denoted $\Delta_{r}$ and $\Delta_{g}$, with $\Delta\in\{\Delta_{r},\Delta_{g}\}$.
+A binary function is defined as:
+$f(\delta,\Delta)=\begin{cases}1&\text{if }\|\delta-\delta^{\prime}\|_{2}\leq\|\delta^{\prime}-\text{NN}_{k}(\delta^{\prime},\Delta)\|_{2}\text{ for at least one }\delta^{\prime}\in\Delta,\\ 0&\text{otherwise,}\end{cases}$
+where $\text{NN}_{k}(\delta^{\prime},\Delta)$ denotes the $k$-th nearest neighbor of $\delta^{\prime}$ in $\Delta$ (we use $k=3$ in all experiments).
+The function $f(\delta_{g},\Delta_{r})$ determines whether a generated sample appears realistic, while $f(\delta_{r},\Delta_{g})$ assesses whether a real sample could be reproduced by the generator.
+Precision and Recall are defined as:
+$\text{Precision}(\Delta_{r},\Delta_{g})$ $=\frac{1}{|\Delta_{g}|}\sum_{\delta_{g}\in\Delta_{g}}f(\delta_{g},\Delta_{r}),$
+$\text{Recall}(\Delta_{r},\Delta_{g})$ $=\frac{1}{|\Delta_{r}|}\sum_{\delta_{r}\in\Delta_{r}}f(\delta_{r},\Delta_{g}).$
+Finally, the F1 score is computed as:
+$F1=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}.$
+We note that Precision, Recall, and F1 range from 0 to 1, with 1 indicating optimal performance.
+5.2 Evaluation procedure
+We train all models on imbalanced datasets, where the class frequency decreases as a function of the class index, and evaluate them on balanced test sets.
+This protocol measures robustness by assessing a model’s ability to generate high-quality samples across all classes, regardless of their frequency during training.
+To evaluate performance across varying complexities, we utilize SVHN (Netzer et al., 2011), CIFAR100 (Krizhevsky, 2009; Cao et al., 2019), and CelebA (Liu et al., 2015) (see Appendix D.1).
+SVHN serves as a tractable baseline—simple enough to ensure convergence, yet rich enough to highlight generative discrepancies.
+CIFAR100 provides a more challenging benchmark with high semantic diversity, particularly testing performance in low-data regimes.
+Finally, CelebA enables a detailed analysis of generative quality and class balance in the presence of natural attributes.
+We construct Long-Tail (LT) variants of SVHN and CIFAR100 to explicitly control the training imbalance.
+Specifically, we induce imbalance via an exponential decay in sample counts after equalizing the initial class sizes.
+The imbalance ratio $\rho$ defines the disparity between the most and least frequent classes, with the sample count $M_{y_{i}}$ for class $y_{i}$ given by:
+$M_{y_{i}}=M\cdot\rho^{-\frac{y_{i}-1}{K-1}},$
+where $M$ denotes the original sample count per class and $y_{i}\in\{1,\dots,K\}$ represents the class index.
+For CelebA, we exploit the inherent imbalance of facial attributes.
+We select Mustache and Young to represent strong ($\rho\approx 25$) and moderate ($\rho\approx 3.5$) imbalance, respectively, alongside the balanced Male and Smiling attributes.
+We deliberately refrain from defining classes via attribute composition; such an approach would lead to a combinatorial expansion of the label space and result in prohibitive sample sparsity per class.
+We employ the FID (Heusel et al., 2017) as the primary global metric to quantify the distribution shift between generated and real samples.
+However, as FID can be biased in low-sample regimes, we complement it with Precision, Recall, and F1-score to provide a granular, per-class evaluation of the proposed model.
+We detail our experimental settings in Appendix D.
+5.3 $\tau$ parameter study
+Figure 1 illustrates the impact of the sampling standard deviation $\tau$ on FID.
+Models based on the Student’s $t$-distribution benefit from higher standard deviation on CIFAR100-LT compared to SVHN-LT.
+Notably, C-$t^{3}$VAE outperforms C-VAE for $\tau\in[0.25;0.55]$ on CIFAR100-LT, $\tau\in[0.19;0.28]$ on SVHN-LT, and for all $\tau$ values on CelebA.
+Additionally, it surpasses the $t^{3}$VAE FID for all $\tau$ values and across all datasets, underlining the importance of equal per-class prior mass.
 (a) SVHN-LT
@@
 Figure 1: FID score as a function of $\tau$ for the $t^{3}$VAE and C-$t^{3}$VAE models.
-Results are for the imbalance ratio $\rho=100$ for the SVHN-LT and CIFAR100-LT, and for the Mustache attribute ($\rho=25$) in the case of the CelebA dataset.
-Other imbalance ratios’ results paint a similar picture and are provided in Appendix F.3.
+Results are for the imbalance ratio $\rho=100$ for the SVHN-LT and CIFAR100-LT, and the Mustache attribute ($\rho=25$) for CelebA.
+Other imbalance ratios’ results paint a similar picture and are provided in Appendix (E.3).
 The horizontal dashed lines is the FID value of the best performing VAE and C-VAE on each dataset and the vertical dashed line is the value of $\tau$ as derived in Eq.
@@
 We note that the used models in these figures have optimized $\beta$ and $\nu$ hyper-parameters.
-5.2.1 $\tau$ parameter study
-Figure 1 illustrates the sampling standard deviation’s $\tau$ impact on the FID metric.
-We observe that models based on the Student’s t-distribution benefit from higher standard deviation on CIFAR100-LT compared to SVHN-LT.
-Specifically, C-$t^{3}$VAE outperforms the C-VAE for $\tau\in[0.25;0.55]$ on CIFAR100-LT and $\tau\in[0.19;0.28]$ on SVHN-LT and for all $\tau$ values on the CelebA dataset.
-Moreover, it surpasses the $t^{3}$VAE models’ FID for all $\tau$ values and across all datasets.
-For both Student’s t-distribution based models, the optimal FID score for the SVHN-LT occurs near the theoretically derived $\tau$ value.
+Moreover, for both Student’s $t$-distribution based models, the optimal FID score for SVHN-LT occurs near the theoretically derived $\tau$ value.
 However, for the more complex CIFAR100-LT dataset, the optimal $\tau$ is higher than the theoretical value $\tau=0.4$.
-We hypothesize that the higher variance reflects the model’s need during training to accommodate the greater complexity and entropy of the dataset, which makes it harder to align the learned approximate posterior with the prior and to produce confident reconstructions.
-For the CelebA dataset, the standard deviation parameter $\tau$ has minimal to no impact on model performance likely due to the lower variability in the dataset’s images as they all represent faces.
-5.2.2 Optimized Model Results Discussion
+The analytically derived $\tau$ provides a principled initialization grounded in divergence geometry.
+Dataset-specific deviations reflect encoder–decoder capacity limits rather than theoretical inconsistency, highlighting interactions between prior heaviness and representation complexity.
+For CelebA, $\tau$ has minimal impact on performance, likely due to the lower variability in the dataset’s images (All faces are visually similar, reducing the impact of latent scale).
+Developing tighter theoretical characterizations of $\tau$ that account for dataset complexity and model capacity remains an interesting direction for future work.
+5.4 Optimized Model Results Discussion
 After optimizing the hyperparameters of the various models tested in this work, we present their generation FID scores in Table 1.
-We provide results for $\beta$ models after optimization and non $\beta$ models ($\beta=1$) to underscore the importance of this parameter on the resulting performance as it was not explored in the $t^{3}$VAE work [10].
+We provide results for optimized $\beta$ models and non-$\beta$ models ($\beta=1$) to underscore the importance of this parameter, which was not explored in the original $t^{3}$VAE work (Kim et al., 2024).
 Table 1: Generation FID results on the SVHN-LT, CIFAR100-LT and CelebA datasets.
@@
 $\beta$-C-$t^{3}$VAE 44.02 42.60 42.01 44.49 103.25 102.99 105.92 112.37 095.82 82.61 081.65 80.08
+From Table 1, $t^{3}$VAE improves over VAE, highlighting the generative advantage of the Student’s $t$-distribution prior over a Gaussian one, in addition to the reconstruction advantage noted in (Kim et al., 2024).
+Optimizing $\beta$ improves $t^{3}$VAE FID over VAE on CIFAR100-LT and CelebA, while remaining competitive on SVHN-LT.
+Qualitative results in Figure 2 show that $t^{3}$VAE produces sharper synthetic images on CelebA than VAE.
 Figure 2: Sample synthetic images from the optimized VAE and $t^{3}$VAE models trained on the CelebA dataset. No class conditioning is possible for these models.
-From Table 1, $t^{3}$VAE consistently improves over the VAE model.
-This underlines the generative advantage of the Student’s t-distribution prior over a Gaussian one, in addition to the reconstruction advantage highlighted by [10].
-Moreover, optimizing the $\beta$ hyper-parameter improves $t^{3}$VAE’s FID over the VAE on the CIFAR100-LT and CelebA, while remaining competitive on SVHN-LT.
-Also, qualitative results of Figure 2 show that the $t^{3}$VAE model compared to the VAE model produces significantly sharper synthetic images on the CelebA dataset.
+For class-conditional models, optimized C-$t^{3}$VAE yields strong FID improvements across imbalanced settings on SVHN-LT and CIFAR100-LT, and on heavily imbalanced CelebA attributes.
+As shown in Table 1, it achieves gains of up to 4, 5, and 10 FID points over $\beta$-$t^{3}$VAE on imbalanced settings of SVHN-LT, CIFAR100-LT, and CelebA, respectively.
+This supports the role of uniform prior mass allocation in high-imbalance regimes.
+Moreover, $\beta$-C-$t^{3}$VAE reduces FID by up to 4 and 15 points over C-VAE on SVHN-LT and CIFAR100-LT, respectively.
+For CelebA, $\beta$-C-$t^{3}$VAE achieves the best results on heavily imbalanced attributes like Mustache, indicating improved generation for underrepresented classes.
+The gain over C-VAE follows from the Student’s $t$-distribution latent prior and its ability to better capture intra-class long-tail structure.
+Qualitative samples of conditional optimized models (Figure 3) show sharper facial features for C-$t^{3}$VAE than C-VAE, notably on the imbalanced Mustache attribute.
+Overall, C-$t^{3}$VAE exhibits the most consistent performance in high-imbalance regimes within the VAE family.
 Figure 3: Sample synthetic images for the optimized C-VAE and C-$t^{3}$VAE trained on specific attributes of the CelebA dataset.
-For class-conditional models, $\beta$ optimized C-$t^{3}$VAE yields strong FID improvements across all imbalance settings, surpassing all baselines.
-As per Table 1, it achieves a gain of up to 4, 5 and 10 FID points over the $\beta$-$t^{3}$VAE model on the imbalanced settings of the SHVN-LT, CIFAR100-LT and CelebA, respectively.
-This underlines the importance of an equal per-class volume in an imbalanced dataset setting.
-Moreover, $\beta$-C-$t^{3}$VAE reduces the FID by up to 4 and 15 points over C-VAE on SVHN-LT and CIFAR100-LT respectively as per Table 1.
-For the CelebA dataset, Table 1 shows that $\beta$-C-$t^{3}$VAE achieves the best results on heavily imbalanced attributes like Mustache, demonstrating better generation for underrepresented groups.
-This gain in performance compared to the C-VAE goes back to the use of Student’s t-distribution latent prior and its ability to better capture intra-class long-tail distributions.
-Additionally, qualitative samples of conditional optimized models (Figure 3) reveal sharper facial features in C-$t^{3}$VAE compared to the C-VAE.
-These results establish it as the most reliable model across imbalance ratios and resolutions.
-In summary, C-$t^{3}$VAE after optimization of all hyper-parameters notably $\beta$, $\nu$ and $\tau$ consistently achieves the lowest FID, validating the use of class-conditional Student’s t-distribution priors and balanced latent sampling.
-Improvements on CIFAR100-LT are especially pronounced, highlighting the model’s robustness to severe imbalance.
-However, for balanced settings, C-$t^{3}$VAE remains competitive.
-5.2.3 Per-class evaluation
+5.5 Per-class evaluation
 In this section, we evaluate the conditional models on a per-class basis.
-However, since FID is biased on small datasets and offers limited insight as a single scalar metric, we rely on Precision, Recall, and F1 generative metrics [^2] Precision quantifies sample quality (sharpness), Recall reflects mode coverage and F1 is their harmornic mean.
-Together they allows us to gain deeper insight into the models’ behavior.
-[26].
-Our results on CelebA are shown in Figure 4, with additional results for SVHN-LT and CIFAR100-LT included in Appendix G.
+Since FID can be biased on small datasets and offers limited insight as a single scalar metric, we rely on Precision, Recall, and F1 metrics (Kynkäänniemi et al., 2019).
+Our results on CelebA are shown in Figure 4, with additional results for SVHN-LT and CIFAR100-LT included in Appendix F.
 (a) Recall
@@
 We note that the imbalance ratio of the Mustache, Young, Male and Smiling factors $\rho$ are $25$, $3.5$, $1.4$ and 11 respectively.
-Figure 5: Fine-grained comparison of C-$t^{3}$VAE and C-VAE models under varying imbalance ratios on SVHN-LT.
-On the CelebA dataset (Figure 4), C-$t^{3}$VAE improves Recall and F1 on the most imbalanced attribute (Mustache), but not on more balanced ones (Male, Smiling) which is an expected behavior.
-However, surprisingly, for the attribute Young (imbalance ratio 3.5), C-VAE performs slightly better this suggests that there is a regime where Gaussian priors may still suffice.
+On CelebA (Figure 4), C-$t^{3}$VAE improves Recall and F1 on the most imbalanced attribute (Mustache), but not on more balanced ones (Male, Smiling), which is expected.
+When class imbalance is moderate (e.g., Young attribute with $\rho=3.5$), C-VAE slightly outperforms C-$t^{3}$VAE, indicating that Gaussian priors suffice when encoder variance and decoder flexibility can compensate for small frequency skew.
+This suggests the presence of a regime where Gaussian priors may suffice.
 To explore this, we vary the imbalance ratio on SVHN-LT from $100$ to $1$ and plot the results in Figure 5.
-This figure shows there is a threshold $\rho\approx 3$ before which C-VAE performs better; beyond that, C-$t^{3}$VAE has the advantage.
-Also, we notice that the more the imbalance ratio increases, the more FID gap becomes more significant and in favor of the C-$t^{3}$VAE.
-For the SVHN-LT dataset (Figures in Appendix G), when comparing the C-$t^{3}$VAE to the C-VAE, the C-$t^{3}$VAE achieves significantly higher Recall and competitive Precision.
-This indicates a better coverage of data modes and maintaining images quality which yields higher F1 scores especially on tail classes.
-For the CIFAR100-LT dataset (Figures in Appendix G), C-VAE often achieves high Precision but zero Recall, indicating mode collapse.
-In contrast, the C-$t^{3}$VAE preserves Recall at the cost of slightly lower Precision, also resulting in improved F1 scores.
-Consequently, on the SVHN-LT CIFAR100-LT and CelebA datasets, C-$t^{3}$VAE sets itself as the most reliable method for fair, high quality image generation.
+This figure shows a threshold around $\rho\approx 5$: below it C-VAE can be better, while above it C-$t^{3}$VAE has the advantage.
+As the imbalance ratio increases, the FID gap grows in favor of C-$t^{3}$VAE.
+For moderate imbalance ($\rho<5$), encoder variance and decoder flexibility can compensate for frequency skew; beyond this regime, posterior–prior mismatch grows nonlinearly, and Gaussian priors increasingly compress minority-class latent regions.
+Heavy-tailed class-conditional priors mitigate this compression by permitting larger dispersion without penalizing minority likelihood, explaining the observed transition.
+Figure 5: C-$t^{3}$VAE versus C-VAE models under varying imbalance ratios on the SVHN-LT dataset.
+For SVHN-LT (Figures in Appendix F), C-$t^{3}$VAE achieves higher Recall and competitive Precision compared to C-VAE, indicating better mode coverage while maintaining image quality, which yields higher F1 scores, especially on tail classes.
+For CIFAR100-LT (Figures in Appendix F), C-VAE often attains high Precision but near-zero Recall, indicating mode collapse.
+In contrast, C-$t^{3}$VAE preserves Recall at the cost of slightly lower Precision, resulting in improved F1 scores.
+Consequently, on SVHN-LT, CIFAR100-LT, and CelebA, C-$t^{3}$VAE is the most reliable method in our study for high-imbalance settings within the VAE family.
 6 Conclusion
-We introduced C-$t^{3}$VAE, a class-conditional generative model that leverages Student’s t-distributions in the latent space with a theoretically derived balanced sampling scheme.
-This approach improves fairness and sample quality under class imbalance.
-Extensive experiments on SVHN-LT, CIFAR100-LT, and CelebA demonstrate that, after the optimization of the $\beta$, $\nu$ and $\tau$ hyper-parameters, C-$t^{3}$VAE consistently outperforms the $t^{3}$VAE and conditional VAE baselines, achieving up to a 15-point FID improvement on highly imbalanced datasets.
-Per-class Precision, Recall, and F1 evaluations confirm improved mode coverage, particularly in tail classes.
-Moreover, we identify $\rho>3$ as the regime where C-$t^{3}$VAE outperforms Gaussian-prior models, with the performance gap widening as $\rho$ increases.
-For $\rho\lesssim 3$, C-$t^{3}$VAE remains competitive with C-VAE.
-Beyond generative performance, C-$t^{3}$VAE contributes to fairness-aware modeling by promoting balanced latent sampling and better mode coverage in underrepresented classes.
-Future work will explore extending C-$t^{3}$VAE to multi-label problems and integration with Latent Diffusion models.
-Acknowledgement
-The authors were granted access to the HPC resources of MesoPSL financed by the Région Île-de-France and the Equip@Meso project (reference ANR-10-EQPX-29-01) of the programme investissements d’avenir supervised by France’s Agence nationale pour la recherche.
-Appendix for
-”Conditional-$t^{3}$VAE: Equitable Latent Space Allocation for Fair Generation”
-Appendix A $\gamma$-power divergence corrected derivation
-In this section, we present our correction of the previously derived form of the $\gamma$-power divergence in [10]. We start from the integrals derived in [10] and keep a similar notation :
-$\int q(x)p(x)^{\gamma}dx$ $=C_{\nu,d}^{\gamma}|\Sigma_{1}|^{-\frac{\gamma}{2}}\left(1+\frac{1}{\nu-2}\operatorname{Tr}(\Sigma_{1}^{-1}\Sigma_{0})+\frac{1}{\nu}(\mu_{0}-\mu_{1})^{\top}\Sigma_{1}^{-1}(\mu_{0}-\mu_{1})\right)$
-$\int q(x)^{1+\gamma}dx$ $=C_{\nu,d}^{\gamma}|\Sigma_{0}|^{-\frac{\gamma}{2}}\left(1+\frac{d}{\nu-2}\right)$
-$\int p(x)^{1+\gamma}dx$ $=C_{\nu,d}^{\gamma}|\Sigma_{1}|^{-\frac{\gamma}{2}}\left(1+\frac{d}{\nu-2}\right)$
-Combining these formulas gives a similar enthropy $\mathcal{H}_{\gamma}$ to the one derived in [10]:
-$\mathcal{H}_{\gamma}(q)=-C_{\nu,d}^{\frac{\gamma}{1+\gamma}}|\Sigma_{0}|^{-\frac{\gamma}{2(1+\gamma)}}\left(1+\frac{d}{\nu-2}\right)^{\frac{1}{1+\gamma}}.$
-However, the cross-entropy $\mathcal{C}_{\gamma}$ takes the following form :
-[…]
-In red, we highlight the main difference to the formula derived in [10]. $\mathcal{H}_{\gamma}(q)$ and $\mathcal{C}_{\gamma}(q,p)$ combine to give :
-$\mathcal{D}_{\gamma}(q\|p)$ $=-\frac{C_{\nu,d}^{\frac{\gamma}{1+\gamma}}}{\gamma}\left(1+\frac{d}{\nu-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg{[}-|\Sigma_{0}|^{-\frac{\gamma}{2(1+\gamma)}}\left(1+\frac{d}{\nu-2}\right)$
-[…]
-Appendix B Priors derivations
+In this work, we introduced C-$t^{3}$VAE.
+This class-conditional generative model mitigates frequency-aligned latent allocation effects observed under imbalance.
+It uses per-class Student’s $t$-distributions in the latent space, paired with a theoretically derived, equal-weight sampling scheme.
+This design allocates uniform prior mass across all classes while capturing heavy-tailed intra-class variations.
+Furthermore, these structural improvements do not sacrifice efficiency.
+The computational complexity of C-$t^{3}$VAE matches that of standard conditional VAEs.
+We evaluated our model on SVHN-LT, CIFAR100-LT, and CelebA.
+After optimizing the hyperparameters $\beta$, $\nu$, and $\tau$, C-$t^{3}$VAE consistently outperforms baseline models in high-imbalance regimes.
+It achieves up to a 15-point FID improvement over $t^{3}$VAE and conditional VAEs.
+Additionally, per-class Precision, Recall, and F1 metrics confirm superior mode coverage for tail classes.
+We also identified a critical imbalance ratio threshold at $\rho\approx 5$.
+For milder imbalances ($\rho\lesssim 5$), Gaussian-based models remain competitive.
+However, as imbalance grows beyond this threshold, C-$t^{3}$VAE demonstrates a significant and widening advantage.
+Future research should explore extending the C-$t^{3}$VAE framework to multi-label settings.
+The MultiFacet VAE model (Falck et al., 2021) provides a promising starting point for handling complex, overlapping annotations.
+Furthermore, our experiments show that the sampling variance $\tau$ heavily impacts image quality.
+This parameter currently requires dataset-specific tuning.
+Therefore, developing tighter theoretical bounds and adaptive latent-space sampling methods remains a crucial next step for high-fidelity VAE generation.
+Acknowledgment
+This work received access to the High-Performance Computing (HPC) resources of MesoPSL, financed by the Region Île-de-France and the Equip@Meso project (reference ANR-10-EQPX-29-01) of the Investissements d’avenir program supervised by the Agence nationale pour la recherche.
+Supplementary Material
+Appendix A Priors derivations
 In this section, we present our derivations of the different prior distributions defining our proposed C-$t^{3}$-VAE model. Starting from the proposed joint distribution :
@@
 $p(z|y)$ $=\int p_{\theta}(x,z|y)dx$
-$=\int C_{\nu,m+n}|\Sigma_{x}|^{-\frac{1}{2}}|\Sigma_{y}|^{-\frac{1}{2}}\Bigg{[}1+\frac{(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})}{\nu}$
-$\quad+\frac{(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))}{\nu}\Bigg{]}^{-\frac{\nu+m+n}{2}}dx$
+$=\int C_{\nu,m+n}|\Sigma_{x}|^{-\frac{1}{2}}|\Sigma_{y}|^{-\frac{1}{2}}\Bigg[1+\frac{(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})}{\nu}+\frac{(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))}{\nu}\Bigg]^{-\frac{\nu+m+n}{2}}dx$
 $=C_{\nu,m+n}|\Sigma_{x}|^{-\frac{1}{2}}|\Sigma_{y}|^{-\frac{1}{2}}\left[1+\frac{1}{\nu}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right]^{-\frac{\nu+m+n}{2}}$
-$\times\int\left(1+\frac{(1+\nu^{-1}m)(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))}{\left(1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right)(\nu+m)}\right)^{-\frac{\nu+m+n}{2}}dx.$
+$\quad\times\int\left(1+\frac{(1+\nu^{-1}m)(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))}{\left(1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right)(\nu+m)}\right)^{-\frac{\nu+m+n}{2}}dx.$
 Given that :
@@
 $=t_{n}\left(x\middle|\mu_{\theta}(z),\frac{\left(1+\nu^{-1}(z-\mu_{y})^{\top}\Sigma_{y}^{-1}(z-\mu_{y})\right)}{(1+\nu^{-1}m)}\Sigma_{x},\nu+m\right).$
-Appendix C Loss function derivation
+Appendix B Loss function derivation
 In this section, we derive the loss function of C-$t^{3}$-VAE.
@@
 $\iint q_{\phi}(x,z|y)p_{\theta}(x,z|y)^{\gamma}dxdz$ $=C_{\nu,m+n}^{\gamma}|\Sigma_{x}|^{-\frac{\gamma}{2}}|\Sigma_{y}|^{-\frac{\gamma}{2}}\,|\Sigma_{y}|^{-\frac{1}{2}}\,\mathbb{E}_{x}\Bigg[1+\frac{1}{\nu}\frac{\nu\operatorname{Tr}\left(\Sigma_{y}^{-1}\Sigma_{\phi}(x)\right)}{\nu+n-2}$
-$\quad+\frac{(\mu(x)-\mu_{y})^{\top}\Sigma_{y}^{-1}(\mu(x)-\mu_{y})}{\nu}$
-$\quad+\frac{\mathbb{E}_{z}\left[(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))\right]}{\nu}\Bigg{]}.$
+$\quad+\frac{(\mu(x)-\mu_{y})^{\top}\Sigma_{y}^{-1}(\mu(x)-\mu_{y})}{\nu}+\frac{\mathbb{E}_{z}\left[(x-\mu_{\theta}(z))^{\top}\Sigma_{x}^{-1}(x-\mu_{\theta}(z))\right]}{\nu}\Bigg].$
 Finally, the third term $\iint q_{(}x,z|y)^{1+\gamma}dxdz$ is
@@
 […]
-Appendix D Sampling distribution variance derivation
+Appendix C Sampling distribution variance derivation
 In this section, we present the derivation of $\tau^{2}$ used in the sampling of $t^{3}$VAE and C-$t^{3}$VAE model.
@@
 First, we simplify the divergence $\mathcal{D}(q\|p^{\star})$ :
-$\mathcal{D}(q\|p^{\star})$ $=-\frac{C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}}{\gamma}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg{[}-\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}$
-$\quad\times\left(1+\frac{m}{\nu+n-2}\right)+|\tau^{2}I|^{-\frac{\gamma}{2}}|\left(1+\nu^{-1}n\right)^{-1}\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}$
-$\quad\times\left(1+\frac{\operatorname{Tr}\left(\tau^{-2}\left(1+\nu^{-1}n\right)^{-1}\Sigma_{\phi}(x)\right)}{\nu+n-2}+\frac{\tau^{-2}}{\nu+n}\|\mu(x)-\mu_{y})\|^{2}\right)\Bigg{]}$
+$\mathcal{D}(q\|p^{\star})$ $=-\frac{C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}}{\gamma}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg[-\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)$
+$\quad+|\tau^{2}I|^{-\frac{\gamma}{2(1+\gamma)}}\times\left(1+\frac{\operatorname{Tr}\left(\tau^{-2}\left(1+\nu^{-1}n\right)^{-1}\Sigma_{\phi}(x)\right)}{\nu+n-2}+\frac{\tau^{-2}}{\nu+n}\|\mu(x)-\mu_{y})\|^{2}\right)\Bigg]$
 Here, we use the fact that $|\alpha A|^{\delta}=\alpha^{\delta n}|A|^{\delta}$ where $n$ is the dimension of the square $A$ matrix.
@@
 After simplification and rearranging we get :
-$\mathcal{D}(q\|p^{\star})$ $=-\frac{1}{\gamma}C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg{[}-\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}$
-$\quad\times\left(1+\frac{m}{\nu+n-2}\right)+\tau^{-\gamma m}\left(1+\nu^{-1}n\right)^{-\frac{\gamma^{2}m}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}$
-$\quad\left(1+\frac{\tau^{-2}\left(1+\nu^{-1}n\right)^{-1}}{\nu+n-2}\operatorname{Tr}\left(\Sigma_{\phi}(x)\right)+\frac{\tau^{-2}}{\nu+n}\|\mu(x)-\mu_{y})\|^{2}\right)\Bigg{]}$
-$=-\frac{1}{\gamma}C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg{[}-\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}$
-$\quad\times\left(1+\frac{m}{\nu+n-2}\right)+\nu^{-1}\tau^{-2-\gamma m}\left(1+\nu^{-1}n\right)^{-\frac{\gamma^{2}m}{2(1+\gamma)}-1}|\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}$
-$\quad\left(\kappa+\frac{\nu}{\nu+n-2}\operatorname{Tr}\left(\Sigma_{\phi}(x)\right)+\|\mu(x)-\mu_{y})\|^{2}\right)\Bigg{]}$
-$=-\frac{1}{\gamma}C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\nu^{-1}\tau^{-2-\gamma m}\left(1+\nu^{-1}n\right)^{-\frac{\gamma^{2}m}{2(1+\gamma)}-1}|\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}$
-$\quad\times\Bigg{[}-\nu\tau^{2+\gamma m}\left(1+\nu^{-1}n\right)^{\frac{\gamma^{2}m+\gamma m}{2(1+\gamma)}+1}|\Sigma_{\phi}(x)|^{-\frac{\gamma^{2}}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)$
-$\quad+\kappa+\frac{\nu}{\nu+n-2}\operatorname{Tr}\left(\Sigma_{\phi}(x)\right)+\|\mu(x)-\mu_{y})\|^{2}\Bigg{]},$
+$\mathcal{D}(q\|p^{\star})$ $=-\frac{1}{\gamma}C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg[-\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)$
+$\quad+\tau^{\frac{-\gamma m}{1+\gamma}}\left(1+\frac{\tau^{-2}\left(1+\nu^{-1}n\right)^{-1}}{\nu+n-2}\operatorname{Tr}\left(\Sigma_{\phi}(x)\right)+\frac{\tau^{-2}}{\nu+n}\|\mu(x)-\mu_{y})\|^{2}\right)\Bigg]$
+$=-\frac{1}{\gamma}C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\Bigg[-\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}}|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)$
+$\quad+\frac{1}{\nu+n}\tau^{-2-\frac{\gamma m}{1+\gamma}}\left(\kappa+\frac{\nu}{\nu+n-2}\operatorname{Tr}\left(\Sigma_{\phi}(x)\right)+\|\mu(x)-\mu_{y})\|^{2}\right)\Bigg]$
+$=-\frac{1}{\gamma}C_{\nu+n,m}^{\frac{\gamma}{1+\gamma}}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\frac{1}{\nu+n}\tau^{-2-\frac{\gamma m}{1+\gamma}}\Bigg[-(\nu+n)\tau^{2+\frac{\gamma m}{1+\gamma}}\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}}$
+$\quad|\Sigma_{\phi}(x)|^{-\frac{\gamma}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)+\kappa+\frac{\nu}{\nu+n-2}\operatorname{Tr}\left(\Sigma_{\phi}(x)\right)+\|\mu(x)-\mu_{y})\|^{2}\Bigg],$
 with:
-$\kappa=\nu\tau^{2}\left(1+\nu^{-1}n\right).$
+$\kappa=\tau^{2}(\nu+n).$
 Then, we match the result to the loss function in Eq. (9) to get :
-$\tau^{2+\gamma m}\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2}+1}|\Sigma_{\phi}(x)|^{-\frac{\gamma^{2}}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)=\frac{C_{1}}{C_{2}}.$
+$\tau^{2+\frac{\gamma m}{1+\gamma}}\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}+1}\left(1+\frac{m}{\nu+n-2}\right)=\frac{C_{1}}{C_{2}}.$
 Moreover, we have :
@@
 Consequently we obtain :
-$\tau^{2+\gamma m}$ $\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2}+1}|\Sigma_{\phi}(x)|^{-\frac{\gamma^{2}}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)$
-$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\;\,=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{-\gamma}{1+\gamma}}_{\nu,n}\left(1+\frac{m}{\nu+n-2}\right)^{\frac{1}{1+\gamma}}\left(1+\frac{m+n}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
-$\tau^{2+\gamma m}$ $=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{-\gamma}{1+\gamma}}_{\nu,n}\left(1+\nu^{-1}n\right)^{\frac{-\gamma m}{2}-1}|\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)^{-1}\left(1+\frac{m}{\nu+n-2}\right)^{\frac{1}{1+\gamma}}$
-$\quad\times\left(1+\frac{m+n}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
-$=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{-\gamma}{1+\gamma}}_{\nu,n}\left(1+\nu^{-1}n\right)^{-\frac{\gamma m}{2}-1}|\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\left(1+\frac{m+n}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
-$=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{-\gamma}{1+\gamma}}_{\nu,n}\left(1+\nu^{-1}n\right)^{-\frac{\gamma m}{2}-1}|\Sigma_{\phi}(x)|^{\frac{\gamma^{2}}{2(1+\gamma)}}\left(\frac{\nu+n-2}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
-$=\left(1+\nu^{-1}n\right)^{-\frac{\gamma m}{2}-1}\left(\sigma^{n}C^{-1}_{\nu,n}|\Sigma_{\phi}(x)|^{\frac{\gamma}{2}}\quad\frac{\nu+n-2}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}.$
+$\tau^{2+\frac{\gamma m}{1+\gamma}}\left(1+\nu^{-1}n\right)^{\frac{\gamma m}{2(1+\gamma)}+1}\left(1+\frac{m}{\nu+n-2}\right)=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{-\gamma}{1+\gamma}}_{\nu,n}\left(1+\frac{m}{\nu+n-2}\right)^{\frac{1}{1+\gamma}}\left(1+\frac{m+n}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
+$\tau^{2+\frac{\gamma m}{1+\gamma}}$ $=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{-\gamma}{1+\gamma}}_{\nu,n}\left(1+\nu^{-1}n\right)^{-\frac{\gamma m}{2(1+\gamma)}-1}\left(1+\frac{m}{\nu+n-2}\right)^{-\frac{\gamma}{1+\gamma}}\left(1+\frac{m+n}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
+$=\sigma^{\frac{n\gamma}{1+\gamma}}C^{\frac{-\gamma}{1+\gamma}}_{\nu,n}\left(1+\nu^{-1}n\right)^{-\frac{\gamma m}{2(1+\gamma)}-1}\left(\frac{\nu+n-2}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}$
+$=\left(1+\nu^{-1}n\right)^{-\frac{\gamma m}{2(1+\gamma)}-1}\left(\sigma^{n}C^{-1}_{\nu,n}\quad\frac{\nu+n-2}{\nu-2}\right)^{\frac{\gamma}{1+\gamma}}.$
 Hence, we get :
-$\tau^{2}$ $=\left(1+\nu^{-1}n\right)^{-\frac{\gamma m+2}{\gamma m+2}}\left(\sigma^{n}C^{-1}_{\nu,n}|\Sigma_{\phi}(x)|^{\frac{\gamma}{2}}\frac{\nu+n-2}{\nu-2}\right)^{\frac{2\gamma}{(1+\gamma)(2+\gamma m)}}$
-$=\left(1+\nu^{-1}n\right)^{-1}\left(\sigma^{n}C^{-1}_{\nu,n}|\Sigma_{\phi}(x)|^{\frac{\gamma}{2}}\frac{\nu+n-2}{\nu-2}\right)^{\frac{2\gamma}{(1+\gamma)(2+\gamma m)}},$
+$\tau^{2}=\left(1+\nu^{-1}n\right)^{-1}\left(\sigma^{n}C^{-1}_{\nu,n}\quad\frac{\nu+n-2}{\nu-2}\right)^{-\frac{2}{\nu+n-2}}.$
 which is the form of $\tau^{2}$ we report in Eq. (8).
-Appendix E Experimental Setup Details
-E.1 Datasets
-We conduct experiments on three datasets notably SVHN-LT [11], CIFAR100-LT [24, 12] and CelebA [13] each chosen to highlight different challenges related to generative modeling under class imbalance and varying visual complexity.
-  • SVHN-LT : This dataset is comprised of colored images of digits from 0 to 9 of size $32\times 32\times 3$.
-It serves as our controlled experimental setting.
-While simple enough for all models to converge, it is rich enough to reflect performance differences.
-However, as this dataset is naturally imbalanced, we balance the number of images across classes to have more control over the imposed imbalance ratio.
+Appendix D Experimental setup
+D.1 Datasets
+We conduct experiments on three datasets notably SVHN-LT (Netzer et al., 2011), CIFAR100-LT (Krizhevsky, 2009; Cao et al., 2019) and CelebA (Liu et al., 2015) each chosen to highlight different challenges related to generative modeling under class imbalance and varying visual complexity.
+  • SVHN-LT : The Street View House Numbers (SVHN) dataset (Netzer et al., 2011) is composed of real-world digit images collected from Google Street View.
+It contains more than 600,000600{,}000 labeled digits (099) with size $32\times 32$, complex backgrounds, and diverse illumination conditions.
+The digits in SVHN are not centered or uniformly scaled, which makes the dataset considerably more challenging.
+Figure 6 shows sample images of this dataset.
+    Figure 6: Sample images from the SVHN dataset (Netzer et al., 2011).
+    However, as this dataset is naturally imbalanced, we balance the number of images across classes to have full control over the imposed imbalance ratio.
 In Table 2 we provide the number of images present in the dataset before balancing.
-    Table 2: The Number of images in the SVHN dataset for the train and test sets before balancing. The value in bold is the one used to balance the dataset.
     Class 0 1 2 3 4 5 6 7 8 9
     Train set 4948 13861 10585 8497 7458 6882 5727 5595 5045 4656
-    Test set 1744 5099 4149 2882 2523 2384 1977 2019 1660 1595
-  • CIFAR100-LT : This balanced dataset comprised of colored images of 100 classes of different natural objects.
-The images are of size $32\times 32\times 3$ presents high variability with 100 fine-grained classes.
-In the most extreme imposed imbalance setting, tail classes may contain as few as five examples.
-Though the images remain low-resolution, this setting stresses the models’ robustness in the face of sparse data and various categories.
-  • CelebA : This dataset enables evaluation in real-world class-imbalance scenarios using attribute labels (e.g., Mustache, Young).
-It also introduces the complexity of higher resolution images ($178\times 218\times 3$), testing the scalability of our models.
-We select four binary attributes (Mustache, Young, Male, Smiling) for evaluation, with test sets balanced by down-sampling the larger class.
-In Table 3, we provide the number of images for each chosen attribute in this dataset.
-    Table 3: The Number of images in the CelebA dataset for the train and test sets for the Mustache, Young, Male and Smiling attributes.
+    Test set 1744 05099 04149 2882 2523 2384 1977 2019 1660 1595
+    Table 2: The Number of images in the SVHN dataset for the train and test sets before balancing. The value in bold is the one used to balance the dataset.
+    For both training and testing, we crop each class to the minimum number of samples available across all classes.
+The only data augmentation applied is a random horizontal flip with 50% probability.
+  • CIFAR100-LT : The CIFAR100 dataset (Krizhevsky, 2009) consists of 60,00060{,}000 color images of size $32\times 32$ pixels, evenly distributed across $100$ object categories.
+Each category contains $600$ images, split into $500$ training samples and $100$ test samples.
+Figure 7 shows sample images after preprocessing.
+    Figure 7: Sample images from the CIFAR100 dataset (Krizhevsky, 2009).
+    We use the dataset in its entirety without class filtering. As with SVHN-LT, we apply a random horizontal flip with 50% probability for data augmentation.
+  • CelebA : The CelebFaces Attributes Dataset (CelebA) (Liu et al., 2015) contains 202,599 color images of celebrity faces at a resolution of $178\times 218$.
+Each image is annotated with 40 binary facial attributes, such as Mustache, Smiling, and Young.
+This dataset exhibits significant variability in pose, expression, and illumination while providing high resolution images.
+We preprocess the images of this dataset by cropping the central region to $160\times 160$ and resizing to $128\times 128$ using bilinear interpolation.
+Figure 8 illustrates sample images after preprocessing.
+Also, we adhere to the original training, validation, and test splits provided by the dataset.
+    Figure 8: Sample images from the CelebA dataset after preprocessing (Liu et al., 2015).
     Mustache No Mustache Young Old Male Female Smiling Not Smiling
     Train set 6642 156128 126788 35982 68261 94509 78080 84690
-    Test set 722 19190 15114 4848 7715 12247 9987 9975
-E.2 Data Preprocessing
-  • SVHN-LT Dataset For both training and testing, we crop each class to the minimum number of samples available across all classes.
-The only data augmentation applied is a random horizontal flip with 50% probability.
-  • CIFAR100-LT Dataset We use the dataset in its entirety without class filtering. As with SVHN-LT, we apply a random horizontal flip with 50% probability for data augmentation.
-  • CelebA Dataset We use the full dataset and apply a center crop of $160\times 160$ to each image and then resize to $128\times 128$. A random horizontal flip with 50% probability is also applied.
-E.3 Model Architecture
+    Test set 0722 019190 015114 04848 07715 12247 09987 09975
+    Table 3: Number of images in the CelebA dataset for the train and test sets for the Mustache, Young, Male and Smiling attributes.
+    Given CelebA’s multi-attribute structure, we select four binary attributes—Mustache, Young, Male, and Smiling—for training and evaluation, treating each attribute and its negation as distinct classes.
+These attributes were chosen due to their varying imbalance ratios, defined as the frequency of the majority class divided by the minority.
+The resulting ratios are 25 for Mustache, 3.5 for Young, 1.4 for Male, and 1 for Smiling.
+This enables the study of imbalance effects in generative modeling.
+We also balance the test sets by downsampling the larger class for each attribute.
+In Table 3 we report the number of images per selected attribute in the dataset.
+D.2 Model Architecture
 Our encoder-decoder models follow a modular block design.
@@
 The decoder has six transposed convolutional layers with channels {512, 512, 256, 128, 64, 32}, followed by a final convolutional layer and Sigmoid activation.
-E.4 Training Details
+D.3 Training Details
 All models are trained using the AdamW optimizer with a learning rate of $10^{-3}$ for 150 epochs. We use a batch size of 64 for SVHN-LT and CIFAR100-LT, and 128 for CelebA.
-Appendix F Hyperparameter Tuning
+Appendix E Hyperparameter Tuning
 We present the hyperparameter tuning process used across all evaluated models. We first optimize $\beta$, then $\nu$, and finally $\tau$, yielding the models’ results reported in Table 1.
-F.1 $\beta$ Optimization
+E.1 $\beta$ Optimization
 We perform a hyperparameter study over $\beta$ for all tested models. Unless otherwise noted, we use the theoretically derived $\tau^{2}$ and set $\nu=10$.
-F.1.1 On the SVHN-LT dataset
+E.1.1 On the SVHN-LT dataset
 As shown in Figure 9, the optimal $\beta$ values for Student’s t models lies in the range $\beta\in[0.4,0.6]$ whereas it lies in the $\beta\in[0.05,0.07]$ range for Gaussian-based models.
@@
 (b) SVHN-LT $\rho$=50
-(a) SVHN-LT $\rho$=10
-(b) SVHN-LT $\rho$=1
+(c) SVHN-LT $\rho$=10
+(d) SVHN-LT $\rho$=1
 Figure 9: Variability of the FID as a function of the $\beta$ hyperparameter for the VAE, C-VAE, $t^{3}$VAE and C-$t^{3}$VAE on the SVHN-LT dataset.
-F.1.2 On the CIFAR100-LT dataset
+E.1.2 On the CIFAR100-LT dataset
 From Figure 10, we observe that Student’s t models obtain the best performance in terms of FID at $\beta=0.2$ for CIFAR100-LT dataset.
@@
 (b) CIFAR100-LT $\rho$=50
-(a) CIFAR100-LT $\rho$=10
-(b) CIFAR100-LT $\rho$=1
+(c) CIFAR100-LT $\rho$=10
+(d) CIFAR100-LT $\rho$=1
 Figure 10: Variability of the FID as a function of the $\beta$ hyperparameter for the VAE, C-VAE, $t^{3}$VAE and C-$t^{3}$VAE on the CIFAR100-LT dataset.
-F.1.3 On the CelebA dataset
-For the CelebA dataset, we focus only on $\beta$ optimization for Student’s t models.
-As for the Gaussian-based models, we set $\beta=0.1$, because as shown in Table 1 the optimization of the $\beta$ hyperparameter has marginal impact on the FID for the CelebA dataset, which is not the case for the SVHN-LT and CIFAR100-LT dataset.
-Hence, we did not deem necessary to perform a hyper-parameter optimization over $\beta$ for the Gaussian based models.
-Nevertheless, from Figure 11 we see that $\beta$ has marginal effect on the Student’s t based models too, likely due to the low intra-class variability.
+E.1.3 On the CelebA dataset
+For CelebA, we optimize $\beta$ exclusively for Student’s $t$ models, setting $\beta=0.1$ for Gaussian variants.
+Table 1 indicates that $\beta$ has minimal impact on CelebA’s FID, unlike the trends seen in SVHN-LT and CIFAR100-LT.
+Therefore, extensive tuning for Gaussian models is omitted.
+Figure 11 suggests that Student’s t models share this robustness to $\beta$ variations, likely driven by the dataset’s limited intra-class variability.
 (a) CelebA - Mustache
 (b) CelebA - Young
-(a) CelebA - Male
-(b) CelebA - Smiling
+(c) CelebA - Male
+(d) CelebA - Smiling
 Figure 11: Variability of the FID as a function of the $\beta$ hyperparameter for the $t^{3}$VAE and C-$t^{3}$VAE on the CelebA dataset.
 The horizontal lines for the VAE and C-VAE models are for the best performing model between $\beta=0.1$ and $\beta=1$.
-F.2 $\nu$ Optimization
-Table 4 shows results from tuning the degrees of freedom parameter $\nu$ in C-$t^{3}$VAE on both SVHN-LT and CIFAR100-LT across all imbalance ratios, using the optimal $\beta$ from the previous study.
-On average, $\nu=10$ performs well, consistent with prior work [10].
-However, performance can be further improved by selecting $\nu$ within the range $[2.5,20]$.
-Still, no major influence of hyper-parameter $\nu$ on the generative FID can be observed similarly to what was observed by [10] for the reconstruction FID.
+E.2 $\nu$ Optimization
+Table 4 presents a sensitivity analysis of the degrees of freedom parameter $\nu$ for the C-$t^{3}$VAE on SVHN-LT and CIFAR100-LT, using the optimal $\beta$ from the previous study.
+Consistent with prior work (Kim et al., 2024), we find that $\nu$ = 10 yields robust performance on average, though slight gains can be achieved by fine-tuning within the range $[2.5,20]$.
+Ultimately, however, the generative FID remains relatively insensitive to variations in $\nu$, corroborating the findings of (Kim et al., 2024) regarding reconstruction quality.
 SVHN-LT CIFAR100-LT
@@
 100 45.97 44.63 43.74 47.52 123.26 122.90 127.42 125.67
-Table 4: Variability of the FID as a function of the standard deviation $\nu$ for the C-$t^{3}$VAE model on the SVHN-LT and CIFAR100-LT datasets.
-F.3 $\tau$ Optimization
+Table 4: Variability of the FID as a function of the standard deviation $\nu$ for the C-$t^{3}$VAE model.
+E.3 $\tau$ Optimization
 In this section, we evaluate the effect of the $\tau$ parameter on the SVHN-LT, CIFAR100-LT and CelebA datasets for all imbalance ratios while setting $\beta$ and $\nu$ to their previously optimized values.
@@
 On CelebA, $\tau$ has minimal impact and the most likely value is $\tau\approx 0.3$.
-This value is between the value observed in for the SVHN-LT and CIFAR100-LT, as is mostly liked to the complexity and the entropy of the dataset where the more complex the data is the higher the value of $\tau$ is required.
+The analytically derived $\tau$ provides a principled initialization grounded in divergence geometry, while dataset-specific deviations reflect encoder–decoder capacity limits rather than theoretical inconsistency, highlighting the interaction between prior heaviness and representation complexity.
 (a) SVHN-LT $\rho$=50
@@
 (f) CIFAR100-LT $\rho$=1
-(a) CelebA Young
-(b) CelebA Male
-(c) CelebA Smiling
+(g) CelebA Young
+(h) CelebA Male
+(i) CelebA Smiling
 Figure 12: Variability of the FID as a function of the standard deviation $\tau^{2}$ for the $t^{3}$VAE and C-$t^{3}$VAE.
@@
 In vertical dashed lines is the theoretically identified value of $\tau$.
-Appendix G Per-Class Evaluation
+Appendix F Per-Class Evaluation
 In this section, we assess the conditional models’ per-class Recall, Precision, and F1 metrics under all imbalance settings and for all tested datasets after optimization of all hyper-parameters.